### v1

# Enumerative Geometry of Quantum Periods

Abstract.
We interpret the $q$-refined theta function $\vartheta_{1}$ of a log Calabi-Yau surface $(\mathbb{P},E)$ as a natural $q$-refinement of the open mirror map, defined by quantum periods of mirror curves for outer Aganagic-Vafa branes on the local Calabi-Yau $K_{\mathbb{P}}$. The series coefficients are all-genus logarithmic two-point invariants, directly extending the relation found in [GRZ]. Yet we find an explicit discrepancy at higher genus in the relation to open Gromov-Witten invariants of the Aganagic-Vafa brane. Using a degeneration argument, we express the difference in terms of relative invariants of an elliptic curve. With $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ the toric blow up of a point, we use the Topological Vertex [AKMV] to show a correspondence between open invariants of $K_{\mathbb{P}}$ and closed invariants of $K_{\widehat{\mathbb{P}}}$ generalizing a variant of [CCLT, LLW] to arbitrary genus and winding. We also equate winding-1, open-BPS invariants with closed Gopakumar-Vafa invariants.
1. Introduction
In [GRZ], the authors found new enumerative interpretations involving periods of the mirror curve of a toric Calabi-Yau threefold, $Y$111 Other enumerative interpretations had previously been found [CLL, CLT, CCLT, CLLT, LLW]. See in particular Theorem 1.1 and Example 4.6 of [LLW] and Example 6.5 (3) of [CCLT]. The conjecture in [GRZ] was partly proved in [You]. The Picard-Fuchs equations determine the periods, and in particular the open mirror map, and its coefficients have two equal interpretations: (1) two-point relative invariants used to define the first theta function, or proper Landau-Ginzburg potential; (2) open invariants of an Aganagic-Vafa brane on $Y.$ In this paper, we extend (1) to higher genus by relating the $q$-refined quantum periods of the quantum mirror curve. We find a discrepancy in (2) at higher genus, and give an explicit expression for the correction by analyzing the degeneration formula. Here, by a $q$-refinement we mean the following.
Definition 1.1.
A q-refinement of $f\in R\llbracket t\rrbracket$ is $g\in R[q^{\pm\frac{1}{2}}]\llbracket t\rrbracket$ such that $g|_{q=1}=f$.
Example 1.2.
For example, when $Y=K_{\mathbb{P}^{2}},$ the periods of (classical) local mirror symmetry are determined by an ODE whose logarithmic solution has holomorphic part
$F(z)=\sum_{k>0}(-1)^{k}\frac{(3k)!}{k(k!)^{3}}z^{k}$
— see, e.g., [GRZ], Section 2.1. The quantum period (see Example 2.15) is
$a(z,q)=-\left(q^{\frac{1}{2}}+q^{-\frac{1}{2}}\right)z-\left(q^{2}+\frac{7}{2}q+6+\frac{7}{2}q^{-1}+q^{-2}\right)z^{2}+\cdots,$
which $q$-refines $-\frac{1}{3}F(-z)=-2z-15z^{2}-\cdots.$ The non-refined period $-\frac{1}{3}F(-z)$ leads to open GW predictions for the AV brane in framing zero, which for genus zero and winding one in low degrees gives $2,5,32,\cdots$. These same numbers were also found in [GRZ] to be two-pointed logarithmic invariants in $\mathbb{P}^{2}$ in genus zero. The same procedure with the quantum period, gives a $q$-refinement of these numbers — specifically
$\left(q^{\frac{1}{2}}+q^{-\frac{1}{2}}\right),\left(q^{2}+q+1+q^{-1}+q^{-2}\right),$
$\left(q^{\frac{9}{2}}+3q^{\frac{7}{2}}+4q^{\frac{5}{2}}+4q^{\frac{3}{2}}+4q^{\frac{1}{2}}+4q^{-\frac{1}{2}}+4q^{-\frac{3}{2}}+4q^{-\frac{5}{2}}+3q^{-\frac{7}{2}}+q^{-\frac{9}{2}}\right),\cdots$
— see Example 2.27. However, the $q$-dependence of the OGW invariants for the AV brane is computed via the topological vertex [AKMV], Equation (9.10) to be
$2,5,(9q+14+9q^{-1}),\cdots,$
quite different.
The aim of this paper is first to prove that the quantum periods computed via the mirror symmetry procedure determine the all-genus logarithmic two-point invariants, then to explain the discrepancy from the OGW invariants.
1.1. Main statements
Here we describe our key findings. The combinatorial Laurent series methods used in §4 work in any dimension, but for the treatment of higher genus invariants with $\lambda$-insertions, we have to restrict to surfaces.
Let $\mathbb{P}$ be a smooth projective Fano surface and let $E$ be a smooth anticanonical divisor, so $(\mathbb{P},E)$ is a log Calabi-Yau pair. Section 1.2 contains our full notation conventions and Section 2 relevant definitions.
1.1.1. Quantum periods
Definition (Definition 2.14).
Let $p(X,Y)$ be the quantum mirror curve defined by $\mathbb{P}$ and write $p=1-f$. We define the quantum A-period by
$a(z,q):={\sf Const}_{X,Y}(\log p(X,Y))=-\sum_{n>0}\frac{1}{n}{\sf Const}_{X,Y}(f^{n}).$
In [ACDKV] the quantum A-period is defined as
$\underline{a}(z,q)={\sf Const}_{X}\left(\frac{\Psi(qX)}{\Psi(X)}\right),$
where the wavefunction $\Psi(X)$ is a solution to the difference equation $p\cdot\Psi=0$ defined by the quantum mirror curve, see §2.5.
Theorem (Theorem 2.23).
The definitions give identical series, $a(z,q)=\underline{a}(z,q)$.
To compute $\underline{a}(z,q)$ from [ACDKV], one can write $Y(X):=\Psi(qX)/\Psi(X)$ as an ansatz $Y(X)=\sum_{k\geq 0}Y_{k}(X)s^{k}$, with $s=q^{\frac{1}{2}}z$ and then solve order-by-order in $s$, see §2.5. Theorem 2.23 states that one can work with the alternative definition $a(z,q)$, which can be computed combinatorially.
Definition (Definition 3.10).
Let $q=e^{i\hbar}$. The $q$-refined regularized GW period of $\mathbb{P}$ is
$\widehat{G}(z,q)=1+\sum_{\beta\in\textup{NE}(\mathbb{P})}\sum_{g\geq 0}(\beta\cdot E)!D_{g,1}(\mathbb{P},\beta)\hbar^{2g}z^{\beta},$
with descendant Gromov-Witten invariants
$D_{g,1}(\mathbb{P},\beta)=\int_{\mathcal{M}_{g,1}(\mathbb{P},\beta)}(-1)^{g}\lambda_{g}\psi^{\beta\cdot{E}-2}\textup{ev}^{\star}[\textup{pt}],$
where $\lambda_{g}=c_{g}(\mathbb{E})$ are the Chern classes of the Hodge bundle.
Theorem (Theorem 3.14).
Let $D$ be defined as in Section 1.2.
$-Da(z,q)=\widehat{G}(z,q).$
The theorem is a $q$-refinement of the mirror description in [CCGGK], Definition 4.9, which states that the regularized quantum period of $\mathbb{P}$ equals the classical period of its mirror potential. Theorem 3.14 is proved using the tropical correspondence for refined descendant invariants of [KSU1].
1.1.2. Natural $q$-refinements of mirror maps
Having a quantum period allows us to use the relation between periods and mirror maps to define $q$-refinements.
Definition (Definition 2.26).
Put $U=xe^{a(-z,q)},$ a relation between two formal parameters $x$ and $U$ via the quantum period. We define $Q^{\beta}=z^{\beta}e^{-(\beta\cdot E)a(-z,q)}$. By Proposition 2.25 there exists an inverse expression $z^{\beta}(Q,q)$. We define the series
$M(Q,q):=e^{-a(-z(Q,q),q)}=\frac{U}{x}=\left(\frac{z^{\beta}(Q,q)}{Q^{\beta}}\right)^{-1/(\beta\cdot E)}$
where the last equality holds for every non-zero class $\beta\in\operatorname{NE}(\mathbb{P})$.
Remark 1.3.
The definition for $M(Q,q)$ is a natural $q$-refinement of the open mirror map $M(Q)$ in the sense that it uses the quantum period of the mirror curve of outer Aganagic-Vafa branes, analogous to how $M(Q)$ is defined. It is not known to us whether $M(Q,q)$ is analytic in a suitable domain. However, it specializes under $q=1$ to the analytic function $M(Q)$.
Definition 3.2,[CPS]).
Let $\vartheta_{1}(y,z,q)_{\infty}$ be the $q$-refined proper Landau-Ginzburg potential (or primitive theta function at infinity) of $(\mathbb{P},{E})$. This is defined as the series enumerating $q$-refined broken lines in a scattering diagram defined by a toric degeneration of $(\mathbb{P},{E})$ in a chamber asymptotically close to $E$. By [GRZ], Theorems 4.14, 4.8 and 5.4, we can equivalently define $\vartheta_{1}(y,z,q)_{\infty}$ as a generating function of $2$-marked logarithmic Gromov-Witten invariants, with $q=e^{i\hbar}$,
$\vartheta_{1}(y,z,q)_{\infty}=y+\sum_{\beta\in\textup{NE}(\mathbb{P})}\sum_{g\geq 0}\frac{1}{\beta\cdot{E}-1}R_{g,(1,\beta\cdot{E}-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}y^{-(\beta\cdot{E}-1)}.$
Theorem (Theorem 4.2).
Under the change of variables $Q^{\beta}=z^{\beta}\cdot(-1/y)^{\beta\cdot E}$ we have
$\vartheta_{1}(y,z,q)_{\infty}=yM(Q,q).$
Corollary (Corollary 4.3).
Let $q=e^{i\hbar}$. Then
$M(Q,q)=1+\sum_{\beta\in\textup{NE}(\mathbb{P})}\sum_{g\geq 0}\frac{1}{\beta\cdot{E}-1}R_{g,(1,\beta\cdot{E}-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}.$
Example 1.4.
For $\mathbb{P}^{2}$, the quantum period is (see Example 2.15),
$a(z,q)=-\left(q^{\frac{1}{2}}+q^{-\frac{1}{2}}\right)z-\left(q^{2}+\frac{7}{2}q+6+\frac{7}{2}q^{-1}+q^{-2}\right)z^{2}+\cdots.$
The natural $q$-refinement of the closed mirror map is
$Q=ze^{-3a(-z,q)}=z-\left(3q^{\frac{1}{2}}+3q^{-\frac{1}{2}}\right)z^{2}+\left(3q^{2}+15q+27+15q^{-1}+3q^{-2}\right)z^{3}+\cdots,$
with inverse
$z(Q,q)=Q+\left(3q^{\frac{1}{2}}+3q^{-\frac{1}{2}}\right)Q^{2}+\left(-3q^{2}+3q+9+3q^{-1}-3q^{-2}\right)Q^{3}+\cdots.$
The definition gives
$\begin{array}[]{l}M(Q,q)=\left(\frac{z(Q,q)}{Q}\right)^{-1/3}=1-\left(q^{\frac{1}{2}}+q^{-\frac{1}{2}}\right)Q+\left(q^{2}+q+1+q^{-1}+q^{-2}\right)Q^{2}\\[5.0pt] \ \ -\left(q^{\frac{9}{2}}+3q^{\frac{7}{2}}+4q^{\frac{5}{2}}+4q^{\frac{3}{2}}+4q^{\frac{1}{2}}+4q^{-\frac{1}{2}}+4q^{-\frac{3}{2}}+4q^{-\frac{5}{2}}+3q^{-\frac{7}{2}}+q^{-\frac{9}{2}}\right)Q^{3}+\cdots.\end{array}$
Theorem 4.2 and Corollary 4.3 extend the results of [GRZ] to the $q$-refined setting. We prove Theorem 4.2 using combinatorial power series results that are proved in §A. These include a Lagrange inversion formula for formal Laurent series and an exponentiation formula for classical periods of these. The latter follows from a relation for partial ordinary Bell polynomials proved in [BGW]. Then Corollary 4.3 follows from the tropical correspondence results in [GRZ]. The combinatorial arguments of §4 work in any dimension. For the q-refined tropical correspondence of §3 we need surfaces.
1.1.3. Open-log correspondence and correction terms
In Section 5, we prove an all genera open-log correspondence between two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ and certain open invariants of $K_{\mathbb{P}}.$ Let $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ be the blow up at a point with exceptional curve $C$, and let $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ be the genus-$g$ Gopakumar-Vafa invariant of $K_{\widehat{\mathbb{P}}}$ in curve class $\pi^{*}\beta-C$ defined by multiple cover formulas [GV1] [GV2]. We first show the following theorem relating $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ and $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$.
Theorem (Theorem 5.2).
Let $(\mathbb{P},E)$ be a smooth log Calabi-Yau pair. Then,
$\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\frac{1}{(\beta\cdot E-1)}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}=\\ \sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\left[(-1)^{\beta\cdot E}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\left(2\sin\frac{\hbar}{2}\right)^{2g-2}Q^{\beta}\right]-\Delta$
where $\Delta$, defined in Equation 5.2, is a discrepancy expressed by the stationary Gromov-Witten theory of $E$ and two-pointed log invariants of $\mathbb{P}(\log E)$.
We prove Theorem 5.2 by using higher genus analogues of arguments in Theorem 1.1, [GRZ]. The discrepancy term $\Delta$ comes from application of the $g\geq 0$ log-local principle of [BFGW] and a blow up formula for log invariants in Corollary 6.6, [GRZ], and is expressed by the stationary Gromov-Witten theory of $E$ and, after fixing a genus-$g$, the two-pointed log invariants $R_{h,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ for $h<g.$
1.1.4. Winding-w open/closed correspondence
In Section 5, we additionally assume that $\mathbb{P}$ is toric and $\pi:\widehat{\mathbb{P}}\to\mathbb{P}$ is a toric blow up, so that $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ are toric Calabi-Yau 3-folds. Let $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be the genus-$g$, winding-$w$, open-BPS invariant of an outer AV-brane $L\subset K_{\mathbb{P}}$ in framing-0, curve class $\beta+w\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$, and in winding profile $\vec{k}$ defined by multiple cover formulas for open Gromov-Witten invariants [MV] (see Section 5 for more details). In Theorem 5.3, we prove an equality between winding-$w$ open partition functions of $K_{\mathbb{P}}$ and a certain closed partition function of $K_{\widehat{\mathbb{P}}}$. Specializing Theorem 5.3 to winding-1, open invariants gives the following correspondence of enumerative invariants.
Theorem (Corollary 5.4).
Suppose $\mathbb{P}$ is a toric Fano surface, $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up at a point with exceptional curve $C$, and let $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ be the toric canonical bundles. Then we have,
$n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)=(-1)^{g+1}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$
In genus-0, Corollary 5.4 is equivalent to Theorem 1.1, [LLW]. In Section 5.6, we provide a computational verification for Corollary 5.4 for $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$ in low degrees and all genera.
By substituting Corollary 5.4 into Theorem 5.2, we show an all genera open-log correspondence in Theorem 5.8 relating two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ to open-BPS invariants $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ for smooth log Calabi-Yau pairs. In genus-00, Theorem 5.8 reduces to the open-log correspondence established in [GRZ]. We give an explicit example of Theorem 5.8 in Example 5.10.
1.2. Notation
Let $\mathbb{P}$ be a smooth projective Fano surface and let $E$ be a smooth anticanonical divisor, so $(\mathbb{P},E)$ is a log Calabi-Yau pair. Let $\textup{NE}(\mathbb{P})$ be the monoid of effective curve classes. For elements of $\mathbb{C}\llbracket\textup{NE}(\mathbb{P})\rrbracket$ we write $z^{\beta}$ or $Q^{\beta}$, depending on the context. We regard $z$ as the complex structure modulus and $Q$ as the symplectic structure modulus. We define an operation $D$ on $\mathbb{C}\llbracket\textup{NE}(\mathbb{P})\rrbracket$ by $Dz^{\beta}=(\beta\cdot E)z^{\beta}$. Note that this operation resembles a log derivative.
The mirror curve is a moduli space for A-branes on the local Calabi-Yau threefold $K_{\mathbb{P}}$, see §2.2. We denote a local coordinate on the mirror curve by $x$. Mirror dual to $x$ is a symplectic area modulus of B-branes, which we denote by $U$.
Let $R$ be a ring and let $f\in R\llbracket y_{1}^{\pm},...,y_{s}^{\pm}\rrbracket$ be a formal Laurent series with coefficients in $R,$ so $f=\sum_{m\in\mathbb{Z}^{s}}f_{m}y^{m}.$ We define ${\sf Const}_{y_{1},\ldots,y_{s}}(f):=f_{0}\in R$ to be the constant term. For a Laurent series $f=\sum_{k\in\mathbb{Z}}a_{k}y^{k}$ and $n\in\mathbb{Z}$, we write $[y^{n}]f=a_{n}$ for the $n$-th coefficient of $f$. In particular, ${\sf Const}_{y}(f)=[y^{0}]f$.
We write $R_{g,(p,q)}(\mathbb{P}(\log E),\beta)$ for the logarithmic Gromov-Witten invariant counting curves on $\mathbb{P}$ of genus-$g$ and class $\beta$ intersecting the divisor $E$ in two points, a prescribed point with multiplicity $p$ and a non-prescribed point with multiplicity $q$. Let $\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$ be a disc class with boundary on an outer Aganagic-Vafa brane $L$. We write $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ for the genus-$g$, winding-$w$, open Gromov-Witten invariant of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}}$ in framing-0 and winding profile $\vec{k}$, and $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be its corresponding open-BPS invariant. We write $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ to be the genus-$g$ local Gromov Witten invariant of $\widehat{\mathbb{P}}$ in curve class $\pi^{*}\beta-wC$ and $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ to be its corresponding Gopakumar-Vafa invariant.
1.3. Acknowledgments
We would like to thank Kwokwai Chan, Melissa Liu, Daniel Krefl, Ezra Getzler and Jörg Teschner for helpful exchanges. T.G. would like to thank the MFO Oberwolfach and the organizers of the workshop Real and Logarithmic Enumerative Geometry for the opportunity to present the results of this project. E.Z. has been supported by NSF grant DMS-2104087. H.R. is supported by the NFR Fripro grant Shape2030.
2. Quantum Mirror Curves, Periods and Wavefunctions
In this section, we give two definitions of quantum periods and show they are equivalent. Local mirror symmetry determines a mirror curve $\Sigma=\{p(X,Y)=0\}\subset(\mathbb{C}^{\times})^{2}$. In the quantum version, the $\mathbb{C}$ coordinates $\log X$ and $\log Y$ are conjugate variables, meaning $YX=qXY$, and the defining equation $p=0$ defines a left ideal for this algebra. A cyclic vector $\Psi$ obeying $p\cdot\Psi=0$ is called a wavefunction. From here there are two routes to defining quantum periods. On the one hand, the classical period of the mirror curve (see [CCGGK]) can be quantized in a natural way. On the other hand, a quantum A-period can be constructed from the wavefunction, following [ACDKV]. In the following subsections, we review these constructions, then prove in Theorem 2.23 that they are equal.
2.1. Fano surfaces
Let $\mathbb{P}$ be a smooth surface with ample anticanonical bundle, i.e., one of the $8$ del Pezzo surfaces given by $\mathbb{P}^{1}\times\mathbb{P}^{1}$ or the blow up of $\mathbb{P}^{2}$ in $0\leq k\leq 6$ points. Let $\mathbb{P}_{\Sigma^{\prime}}$ be a Gorenstein toric Fano surface to which $\mathbb{P}$ admits a $\mathbb{Q}$-Gorenstein degeneration. If $\mathbb{P}$ is toric, we can simply take $\mathbb{P}_{\Sigma^{\prime}}=\mathbb{P}$. Let $\Delta^{\star}$ be the reflexive polygon given by the convex hull of the primitive integral ray generators of $\Sigma^{\prime}$. The ray generators are exactly the vertices of $\Delta^{\star}$. Any mutation of $\Delta^{\star}$ gives another Gorenstein toric Fano surface to which $\mathbb{P}$ admits a $\mathbb{Q}$-Gorenstein degeneration [Ilt]. Let $\Sigma$ be the subdivision of $\Sigma^{\prime}$ whose ray generators are all integral points on the boundary of $\Delta^{\star}$, not just the vertices. Then $\mathbb{P}_{\Sigma}$ is a smooth semi-Fano toric surface to which $\mathbb{P}$ admits a $\mathbb{Q}$-Gorenstein degeneration. Again, if $\mathbb{P}$ is toric we can take $\mathbb{P}_{\Sigma}=\mathbb{P}$. There are exactly 16 possibilities for $\mathbb{P}_{\Sigma^{\prime}}$, hence for $\mathbb{P}_{\Sigma}$, corresponding to the 16 reflexive polygons. See §3.2 and [CPS] for more details and [Gra1], Figure 1.3, for a complete list of cases.
2.2. Definition of the mirror curve
The mirror of the canonical bundle of a Fano surface $K_{\mathbb{P}}$ is a threefold $Z\subset\mathbb{C}^{2}\times(\mathbb{C}^{\times})^{2}$ described by an equation $uv=p(x,y)$. The Riemann surface $\{p=0\}\subset(\mathbb{C}^{\times})^{2}$ is called the (classical) mirror curve. It is the moduli space of holomorphic B-branes supported on $\{u=0\}\cong\mathbb{C}$, mirror to an Aganagic-Vafa A-brane $L\subset K_{\mathbb{P}}$.
We view the mirror curve as being fibered over Spec [$\text{Spec }\mathbb{C}[\textup{NE}(\mathbb{P})]$ via the closed moduli $z^{\beta}$, and we write $p(x,y,z)$. Write $p(x,y,z)=1-f(x,y,z)$. We will give an explicit definition of the mirror curve by specifying $f$.
Linear relations between the ray generators of $\Sigma$ correspond to curve classes of $\mathbb{P}_{\Sigma}$. Choose a cyclic numbering of the rays $\rho_{1},\ldots,\rho_{r+2}$. Let $m_{k}=(a_{k},b_{k})$ be the primitive integral generator of $\rho_{k}$, i.e., $\rho_{k}=\mathbb{R}_{\geq 0}(a_{k},b_{k})$ with $a_{k},b_{k}\in\mathbb{Z}$ and $\gcd(a_{k},b_{k})=1$. A set of independent linear relations
$\sum_{i=1}^{r+2}Q_{ik}m_{k}=0,\quad j=1,\ldots,r,$
can be packed into the charge matrix $Q=(Q_{ik})_{i,k}$. Using Gauß elimination we can transform the first $r$ columns of $Q$ into the identity matrix, so that $Q_{ik}=\delta_{ik}$ for $i,k\leq r$. Since $\mathbb{P}_{\Sigma}$ is smooth, the ray generators of $\rho_{r+1}$ and $\rho_{r+2}$ span the lattice, so all entries of $Q$ are integers. The rows of $Q$ define generators $\beta_{1},\ldots,\beta_{r}$ of the group $H_{2}(\mathbb{P}_{\Sigma},\mathbb{Z})$ by intersections with toric divisors
$D_{\rho_{i}}\cdot\beta_{k}=Q_{ik}.$
If $\Delta^{\star}$ has only edges of lattice length $1$, then $\mathbb{P}_{\Sigma^{\prime}}=\mathbb{P}_{\Sigma}$ is a smooth toric Fano surface, hence isomorphic to $\mathbb{P}^{1}\times\mathbb{P}^{1}$ or the blow up of $\mathbb{P}^{2}$ in $0\leq k\leq 3$ points. In this case, we define $f$ as the Hori-Vafa potential
$f(x,y,z)=\sum_{k=1}^{r+2}z^{\beta_{k}}x^{a_{k}}y^{b_{k}}.$
Now consider the case when $\Delta^{\star}$ has an edge $E$ of length $\ell>1$. Let $(a_{k},b_{k})$ be a lattice point on $E$. Let $m_{k},n_{k}\geq 0$ be the lattice distances of $(a_{k},b_{k})$ from the two vertices of $E$. (If $(a_{k},b_{k})$ is a vertex, then it is contained in two different edges, but the following construction will not depend on this choice.) For a lattice point on $E$ indexed by $k$, let $\tilde{\beta}_{k}$ be the curve class defined by the charge vector $(0,\ldots,0,1,-2,1,0,\ldots,0)$, with the entry $-2$ at the $k$-th position, i.e., by
$D_{\rho_{k}}\cdot\tilde{\beta}_{k}=-2,\quad D_{\rho_{k\pm 1}}\cdot\tilde{\beta}_{k}=1,\quad D_{\rho_{i}}\cdot\tilde{\beta}_{k}=0\ \forall i\notin\{k-1,k,k+1\}.$
Here we consider the indices modulo $r+2$. For an integer $l\geq 0$ and ordered tuples of integers $0\leq i_{1}<\ldots<i_{l}\leq m$ and $0\leq j_{1}<\ldots<j_{l}\leq n$ define
$\beta_{k,(i_{1},\ldots,i_{l}),(j_{1},\ldots,j_{l})}=\beta_{k}+\sum_{p=1}^{l}\sum_{q=0}^{i_{p}-1}\tilde{\beta}_{k-q}+\sum_{p=1}^{l}\sum_{q=0}^{j_{p}-1}\tilde{\beta}_{k+q}.$
Note that if $l=0$, then this is equal to $\beta_{k}$.
Definition 2.1.
Define the toric potential of $\mathbb{P}_{\Sigma}$, resp. $\mathbb{P}$, by
$f(x,y,z)=\sum_{k=1}^{r+2}\left(\sum_{l=0}^{\min(m_{k},n_{k})}\sum_{\begin{subarray}{c}0<i_{1}<\ldots<i_{l}\leq m_{k}\\ 0<j_{1}<\ldots<j_{l}\leq n_{k}\end{subarray}}z^{\beta_{k,(i_{1},\ldots,i_{l}),(j_{1},\ldots,j_{l})}}\right)x^{a_{k}}y^{b_{k}}.$
Remark 2.2.
If $(a_{k},b_{k})$ is a vertex of $\Delta^{\star}$, then $m_{k}=n_{k}=0$ and the coefficient of $x^{a_{k}}y^{b_{k}}$ is $z^{\beta_{k}}$. In particular, it is independent of the choice of edge on which $(a_{k},b_{k})$ lies. Moreover, this shows that $f$ is a generalization of the Hori-Vafa potential.
Lemma 2.3.
The number of summands in the $x^{a_{k}}y^{b_{k}}$-coefficient of $f(x,y,z)$ is $\binom{\ell}{m_{k}}=\binom{\ell}{n_{k}}$.
Proof.
Note that $m_{k}+n_{k}=\ell$, so $\binom{\ell}{m_{k}}=\binom{\ell}{n_{k}}$. Without loss of generality assume $n_{k}\leq m_{k}$. The number of terms in the $x^{a_{k}}y^{b_{k}}$-coefficient of $f(x,y,z)$ is given by
$\sum_{l=0}^{n_{k}}\sum_{\begin{subarray}{c}0<i_{1}<\ldots<i_{l}\leq m_{k}\\ 0<j_{1}<\ldots<j_{l}\leq n_{k}\end{subarray}}1=\sum_{l=0}^{n_{k}}\binom{m_{k}}{l}\binom{n_{k}}{l}.$
The Vandermonde identity states that for all $m,n,r\in\mathbb{N}$ one has
$\binom{m+n}{r}=\sum_{l=0}^{r}\binom{m}{l}\binom{n}{r-l}.$
For $m=m_{k}$, $n=n_{k}$ and $r=n_{k}$ and with $\ell=m_{k}+n_{k}$ the above equation becomes
$\binom{\ell}{m_{k}}=\sum_{l=0}^{n_{k}}\binom{m_{k}}{l}\binom{n_{k}}{n_{k}-l}=\sum_{l=0}^{n_{k}}\binom{m_{k}}{l}\binom{n_{k}}{l}.$
This equation agrees with the above expression for the number of terms. ∎
Remark 2.4.
  1. (1)
    We will see in Lemma 3.6 that $f$ agrees with the primitive theta function $\vartheta_{1}$ in the central chamber of the scattering diagram defined by $\mathbb{P}_{\Sigma}$.
  2. (2)
    The relation obtained from specializing the closed moduli $z^{\beta}$ by pairing with the polarization was already known by [CPS], Theorem 8.2, which amounts to saying that $\binom{\ell}{m_{k}}=\binom{\ell}{n_{k}}$ counts the summands in the $x^{a_{k}}y^{b_{k}}$-coefficient of $\vartheta_{1}$. Lemma 3.6 extends the agreement in part (1) of the remark by giving the correct closed moduli.
  3. (3)
    Moreover, without closed moduli it was shown in [CKPT], Proposition 3.9, that $f$ is the unique maximally mutable Laurent polynomial with Newton polytope $\Delta^{\star}$. It would be interesting to study mutations of $f$ respecting the closed moduli.
  4. (4)
    For non-Fano varieties, the potential $f$ is defined to coincide with the theta function $\vartheta_{1}$ in the central chamber. Compared to Definition 2.1 the potential will receive corrections from internal scattering, see [BGL] and [CPS], §8.2.
Convention 2.5.
By a lattice isomorphism, we can assume that $\Delta^{\star}$ contains the points $(1,0)$ and $(0,1)$ and contains no integral points with second coordinate $>1$. These conventions are satisfied by the cases in [Gra1], Figure 1.3, after changing variables $x\leftrightarrow y$, apart from case (3a) corresponding to a cubic surface, where we have to additionally reflect the picture along the $x$-axis.
From now on we assume Convention 2.5. If $\rho_{j}$ has primitive integral generator $(a_{j},b_{j})$, then the $j$-th row of $Q$ becomes $(0,\ldots,0,1,0,\ldots,0,-a_{j},-b_{j})$, where the entry $1$ is at the $j$-th position. This row corresponds to a curve class $\beta_{j}$ defined by
$D_{\rho_{r+1}}\cdot\beta_{j}=-a_{j},\quad D_{\rho_{r+2}}\cdot\beta_{j}=-b_{j},\quad D_{\rho_{j}}\cdot\beta_{j}=1,\quad D_{\rho_{i}}\cdot\beta_{j}=0\ \forall i\notin\{j,r+1,r+2\}.$
Example 2.6.
Consider $\mathbb{P}^{2}$. A fan $\Sigma$ for $\mathbb{P}^{2}$ is generated by $(-1,-1)$, $(1,0)$ and $(0,1)$. All ray generators are vertices of $\Delta^{\star}=\text{Conv}\{(1,0),(0,1),(-1,-1)\}$. The construction above gives
$f=x+y+\frac{z}{xy},$
where $z$ is the unique closed modulus corresponding to the class of a line. All rays $\rho^{\prime}$ of the fan correspond to toric divisors $D_{\rho^{\prime}}$ whose class is the line class. The ray $\rho=\mathbb{R}_{\geq 0}(-1,-1)$ corresponds to the curve class $\beta_{\rho}$ defined by $\beta_{\rho}\cdot D_{\rho^{\prime}}=1$ for all rays $\rho^{\prime}$, which is equal to the class of a line.
Example 2.7.
Consider $\mathbb{P}^{1}\times\mathbb{P}^{1}$. A fan $\Sigma$ for $\mathbb{P}^{1}\times\mathbb{P}^{1}$ is generated by $(-1,0)$, $(0,-1)$, $(1,0)$ and $(0,1)$. Again, all generators are vertices of $\Delta^{\star}$. The construction above gives
$f=x+y+\frac{z^{\beta_{1}}}{x}+\frac{z^{\beta_{2}}}{y},$
where $\beta_{1}$ and $\beta_{2}$ are the curve classes of the two $\mathbb{P}^{1}$ factors. The class of the toric divisors $D_{\mathbb{R}_{\geq 0}(1,0)}$ and $D_{\mathbb{R}_{\geq 0}(-1,0)}$ is equal to $\beta_{1}$ while the class of $D_{\mathbb{R}_{\geq 0}(0,1)}$ and $D_{\mathbb{R}_{\geq 0}(0,-1)}$ is equal to $\beta_{2}$.
Now, let’s compute the curve classes $\beta_{\rho}$ (note that $\beta_{\rho}$ is not the class of the toric divisor $D_{\rho}$). For $\rho=\mathbb{R}_{\geq 0}(-1,0)$ the corresponding curve class $\beta_{\rho}$ is defined by $\beta_{\rho}\cdot\beta_{1}=1$ and $\beta_{\rho}\cdot\beta_{2}=0$, hence is equal to the class $\beta_{2}$. Analogously, $\beta_{\rho}$ for $\rho=\mathbb{R}_{\geq 0}(0,-1)$ is equal to the class $\beta_{1}$.
Example 2.8.
Again consider $\mathbb{P}^{1}\times\mathbb{P}^{1}$, but this time with a $\mathbb{Q}$-Gorenstein degeneration to the Hirzebruch surface $\mathbb{P}_{\Sigma}=\mathbb{F}_{2}$. The fan $\Sigma$ is generated by $(-1,0)$, $(-2,-1)$, $(1,0)$ and $(0,1)$. The construction above gives
$f=x+y+\frac{z^{F}+z^{F+E}}{x}+\frac{z^{2F+E}}{x^{2}y},$
where $F$ is the class of a fiber $E$ is the class of the negative section. This potential $f$ agrees with the one found in [Aur], Proposition 3.1. See also [BGL], §13.2. Together, $F$ and $E$ generate $\textup{NE}(\mathbb{F}_{2})$, with intersection relations $F^{2}=0$, $E^{2}=-2$ and $F\cdot E=1$. Under the deformation from $\mathbb{F}_{2}$ to $\mathbb{F}_{0}=\mathbb{P}^{1}\times\mathbb{P}^{1}$, the classes are mapped as $F\mapsto\beta_{2}$ and $E\mapsto\beta_{1}-\beta_{2}$.
The toric divisors of $(-1,0)$, $(-2,-1)$, $(1,0)$, $(0,1)$ are given by $E,F,2F+E,F$, in this order. For $\rho_{2}=\mathbb{R}_{\geq 0}(-2,-1)$ the class $\beta_{2}$ is defined by $\beta_{2}\cdot(2F+E)=2$, $\beta_{2}\cdot F=1$ and $\beta_{2}\cdot E=0$. Hence, $\beta_{2}=2F+E$. For $\rho_{1}=\mathbb{R}_{\geq 0}(-1,0)$, the class $\beta_{1}$ is defined by $\beta_{1}\cdot(2F+E)=1$, $\beta_{1}\cdot F=0$ and $\beta_{1}\cdot E=1$. Hence, $\beta_{1}=F$. The point $(-1,0)$ is not a vertex of $\Delta^{\star}$. The class $\tilde{\beta}_{1}$ is defined by $\tilde{\beta}_{1}\cdot E=-2$ and $\tilde{\beta}_{1}\cdot F=1$. Hence, $\tilde{\beta}_{1}=E$ and $\beta_{1,(0),(0)}=\beta_{1}+\tilde{\beta}_{1}=F+E$. Then the $x^{-1}$-coefficient of $f(x,y,z)$ is given by $z^{\beta_{1}}+z^{\beta_{1,(0),(0)}}=z^{F}+z^{F+E}$ as claimed.
2.3. Quantum Mirror Curves and Quantum Periods
The quantum mirror curve arises from topological strings, in which $\log x$ and $\log y$ are conjugate variables that get quantized as $\log X$ and $\log Y$ obeying $[\log Y,\log X]=i\hbar$ ($\hbar$ is also called variously $g_{s}$ and $\lambda$, the string coupling constant). So $X$ and $Y$ generate a quantum torus $YX=qXY,$ with $q=e^{i\hbar}.$
Definition 2.9.
By quantum mirror curve we mean the ideal in this quantum torus defined by $p(X,Y,z)=1-f(X,Y,z)$, for $f$ as in Definition 2.1.
Remark 2.10.
We often suppress the $z$-dependence and write $p(X,Y)=1-f(X,Y)$ as elements in the quantum torus with coefficients in $\mathbb{C}[\textup{NE}(\mathbb{P})]$.
Example 2.11.
The quantum mirror curve for $K_{\mathbb{P}^{2}}$ is defined by
$p\;=1-X-Y-q^{\frac{1}{2}}zX^{-1}Y^{-1}.$
The classical mirror curve is recovered in the limit $\hbar\to 0,$ i.e., $q\to 1.$
Remark 2.12.
A canonical choice of ordering of $p$ can be made as follows. The non-constant monomials in $p$ are of the form $z^{m},$ where $m\in\mathbb{Z}^{2}$ is a ray vector of the fan $\Sigma$ of $\mathbb{P}.$ These terms are quantized as $\hat{z}^{m},$ obeying the commutation relations $\hat{z}^{m}\hat{z}^{m^{\prime}}=q^{-\frac{1}{2}\det(m|m^{\prime})}\hat{z}^{m+m^{\prime}}.$ We take $X=\hat{z}^{(1,0)}$ and $Y=\hat{z}^{(0,1)}.$ Note $\hat{z}^{(-1,-1)}=q^{\frac{1}{2}}\hat{z}^{(-1,0)}\hat{z}^{(0,-1)}=q^{\frac{1}{2}}X^{-1}Y^{-1}$ explains the last term in the quantum mirror curve of $\mathbb{P}^{2}$.
Let $\mathcal{D}$ be the algebra of the quantum torus $YX=qXY$ and let $\mathcal{I}=\mathcal{D}{p}$ be the left ideal defined by ${p}.$ Note that $\mathcal{V}:=\mathbb{C}[\textup{NE}(\mathbb{P})][q^{\pm\frac{1}{2}}][X^{\pm 1}]$ defines a standard representation of $\mathcal{D}$ by $(X\cdot f)(X)=Xf(X)$ and $(Y\cdot f)(X)=f(qX).$ This is the induced representation of $\mathcal{D}$ defined by the representation of the subalgebra $\mathcal{V}\subset\mathcal{D}$ given by $\mathcal{V}$ itself. A wavefunction $\Psi$ for the ideal $\mathcal{I}$ is an element of $\mathcal{V}$ which is a cyclic vector for the $\mathcal{D}$-module $\mathcal{D}/\mathcal{D}{p},$ in plain terms a solution to the difference equation $p\cdot\Psi=0.$
Remark 2.13.
To leading order in $\hbar$, $\Psi\sim e^{W/\hbar}$ where $W$ is the potential of the (classical) mirror curve: $p(x,y)=0,$ i.e.,  $\log y=\partial_{\log x}W.$
2.4. The quantum A-period
According to [CKYZ], §6, one (classical) period to consider is the integral of $\log(p)$ over the torus $|x|=|y|=1.$ We call this the A-period. Note that with $p=-f(x,y,z)$ we have $\log(p)=-\sum_{n\geq 1}\tfrac{1}{n}f^{n}$, and each term $f^{n}$ is a Laurent polynomial in $x$ and $y$. Recall from §1.2 that for such Laurent series we write ${\sf Const}(F)$ for the constant term — then the integral picks up the constant term ${\sf Const}(\log(p))=-\sum_{n\geq 1}\tfrac{1}{n}{{\sf Const}}(f^{n})$. We use this fact to generalize the definition of period to the non-commutative case, with non-commuting variables $X$ and $Y$ obeying $YX=qXY.$
Definition 2.14.
The quantum A-period is
$a(z,q):={\sf Const}_{X,Y}(\log p(X,Y))=-\sum_{n>0}\frac{1}{n}{\sf Const}_{X,Y}(f^{n}).$
Note that when taking the powers $f^{n}$, we have to apply the commutation relation $YX=qXY$ repeatedly, leading to different powers of $q^{\tfrac{1}{2}}$.
Example 2.15.
In the case of $\mathbb{P}=\mathbb{P}^{2}$, we have $p=1-f$ with $f=X+Y+q^{\frac{1}{2}}zX^{-1}Y^{-1}$. Then the constant terms of $f^{3}$ come the $3!$ ways of ordering the product $X\cdot Y\cdot q^{\frac{1}{2}}zX^{-1}Y^{-1}$. Using $XY=q^{-1}YX$ and $YX^{-1}=q^{-1}X^{-1}y,$ the result is seen to be $q^{\frac{1}{2}}(3+3q^{-1}),$ This contributes to the coefficient of $z$ in the quantum period, which we call $a_{1},$ giving $a_{1}(q)=-q^{\frac{1}{2}}-q^{-\frac{1}{2}}$. The next terms are computed in a similar way, giving
$\begin{array}[]{l}a(z,q)=\left({-q^{\frac{1}{2}}-q^{-\frac{1}{2}}}\right)z+\left(-q^{2}-\frac{7}{2}q-6-\frac{7}{2}q^{-1}-q^{-2}\right)z^{2}+\\[5.0pt] \ \ \left(-q^{\frac{9}{2}}-3q^{\frac{7}{2}}-12q^{\frac{5}{2}}-\frac{88}{3}q^{\frac{3}{2}}-48q^{\frac{1}{2}}-48q^{-\frac{1}{2}}-\frac{88}{3}q^{-\frac{3}{2}}-12q^{-\frac{5}{2}}-3q^{-\frac{7}{2}}-{q^{-\frac{9}{2}}}\right)z^{3}+\ldots\end{array}$
Remark 2.16.
The coefficient of $z^{n}$ above is a Laurent polynomial in $q$ which refines the multinomial coefficient $\binom{3n}{n,n,n}$. The coefficient of $q^{k}$ encodes the number of closed paths of length $3n$ in the 2d triangular lattice enclosing a fixed signed area, where only three directions of motion are allowed. For $\mathbb{P}^{1}\times\mathbb{P}^{1}$ we’d be counting closed paths of length $2n$ in the rectangular lattice enclosing a fixed signed area, with steps allowed in any direction , giving a $q$-analogue of $\binom{2n}{n}^{2}$. These $q$-analogues seem to be hard to compute [PS].
2.5. Wavefunctions and the A-period
An important insight of Aganagic-Vafa was that the superpotential $W$ is the Abel-Jacobi map, i.e., the integral of a one-form on the mirror curve: $W=\int\log Yd\log X$ (or equivalently $\log Y=\partial_{\log X}W,$ as in Remark 2.13 above).
It is shown in [ACDKV], Section 2.2 that the one-form $\log Yd\log X$ acquires a pole at the point of brane insertion with residue $\hbar$ (recall $\hbar=g_{s}=\lambda$ is the string coupling) – equivalently $W$ changes by $\hbar$ times a period.222 Quantum mirror curves play a central role in the Topological Strings/Spectral Theory duality of [GHM], where the spectrum of the operator defining the mirror curve is studied. Further period computations can be found in many other works — see, e.g., [DH, HMMO, HK, HW, DL], for example.
By Remark 2.13, since $\Psi\sim e^{W/\hbar}$, then the wavefunction changes by $e^{\textup{period}}$ to leading order. This reasoning leads [ACDKV] to define the quantum A-period as the constant term of $\log(\Psi(x+\hbar)/\Psi(x))$ with respect to the variable $X$:
Definition 2.17.
Define
$\underline{a}(z,q)={\sf Const}_{X}\left(\log\frac{\Psi(qX)}{\Psi(X)}\right).$
We will show in Theorem 2.23 that $\underline{a}(z,q)$ agrees with the quantum A-period of Definition 2.14.
Definition 2.18.
Consider the rings $\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket\textup{NE}(\mathbb{P})\rrbracket$ and $\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]\llbracket\textup{NE}(\mathbb{P})\rrbracket$ with grading given by $\deg(z^{\beta})=\beta\cdot E$ and denote the pieces of degree at least one by $(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$ and $(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$, respectively.
Proposition 2.19.
If $p(X,Y)$ follows the Convention 2.5, the equation $p(X,Y)=0$ has a unique solution $Y=Y(X)$ in $1+X\mathbb{C}[X]+(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$.
Proof.
By Convention 2.5, $f(X,Y)$ contains the monomials $X$ and $Y$. Then the equation $p(X,Y)=0$ modulo $z$ becomes $01-X-Y=$ 0, since all other terms have degree at least one with respect to the grading $\deg(z^{\beta})=\beta\cdot E$. We write as an ansatz $Y(X)=1-X+\sum_{\beta\in\textup{NE}(\mathbb{P})}Y_{\beta}(X)z^{\beta}$. Then we can solve for the terms $Y_{\beta}(X)$ case-by-case in $\beta$ with increasing degree $\deg(z^{\beta})=\beta\cdot E$, such that each $Y_{\beta}(X)$ is an element of $\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]$ of degree at least one. ∎
See §2.6 for the computation of $Y(X)$ for $(\mathbb{P}^{2},E)$. We make no claims about the existence of other solutions.
Proposition 2.20.
Let $Y(X)$ be a solution to $p(X,Y)=0$ and let $\Psi(X)$ be the wavefunction defined by $p\cdot\Psi=0$. Then
(2.1) $Y(X)=\frac{\Psi(qX)}{\Psi(X)},$
Proof.
Let $Y(X)$ be a solution to $p(X,Y(X))=0$ as in Proposition 2.19. Then $p(X,Y(X))\Psi(X)=0$. Hence, the action of $p$ on $\Psi(X)$ is tautologically given by multiplication with $p(X,Y(X))$ and the action of $Y$ on $\Psi(X)$ is given by multiplication with $Y(X)$. Since $Y\cdot\Psi(X)=\Psi(qX)$, this shows $Y(X)=\Psi(qX)/\Psi(X)$. ∎
Lemma 2.21.
Let $Y(X)$ be an element of $1+X\mathbb{C}[X]+(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$. Then $\log Y(X)\in X\mathbb{C}[X]+(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$ and ${\sf Const}_{X}(\log Y(X))\in(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$ are well-defined.
Proof.
We write $Y(X)=1+Y_{1}(X)+Y_{2}(X)$ with $Y_{1}(X)\in X\mathbb{C}[X]$ and $Y_{2}(X)\in(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$. Then
$\log Y(X)=\sum_{k>0}\frac{(-1)^{k}}{k}(Y_{1}(X)+Y_{2}(X))^{k}.$
For all $k,l>0$ we have $Y_{1}(X)^{k}\in X\mathbb{C}[X]$, $Y_{1}(X)^{k}Y_{2}(X)^{l}\in(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$ and $Y_{2}(X)^{k}\in(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$. Taking the formal limit $k,l\rightarrow\infty$ changes $X\mathbb{C}[X]$ to $X\mathbb{C}\llbracket X\rrbracket$ and doesn’t affect $(\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket X\rrbracket[X^{-1}]\llbracket\textup{NE}(\mathbb{P})\rrbracket)_{>0}$. ∎
Corollary 2.22.
We have
$\underline{a}(z,q)={\sf Const}_{X}(\log Y(X)).$
Proof.
The right hand side is well-defined by Lemma 2.21. Then the statement follows directly from Definition 2.17 and Proposition 2.20. ∎
2.6. Computing the quantum A-Period
Here we explain the method of computing the quantum A-period à la [ACDKV], using the example $(\mathbb{P}^{2},E)$. [^3] We thank Daniel Krefl for explaining the calculation in Section 7.3 of [ACDKV]. We have
(2.2) $1-X-Y(X)-\frac{s}{XY(q^{-1}X)}=0,$
where we have put $s=q^{\frac{1}{2}}z$. We will later specialize $s$ in our solution to $q^{\frac{1}{2}}z$.
We write as an ansatz $Y(X)=\sum_{k\geq 0}Y_{k}(X)s^{k},$ then solve order-by-order in $s$. Taking the equation modulo $s$, we get $Y_{0}=1-X.$ Taking (2.2) modulo $s^{2}$ and inserting $Y_{0}$ gives the equation
$1-X-(1-X)-Y_{1}s-\frac{s}{X(1-q^{-1}X+...)}=0,$
so $Y_{1}=-\frac{1}{X(1-q^{-1}X)}.$
Similarly, taking (2.2) modulo $s^{3}$ and inserting $Y_{0}$ and $Y_{1}$ gives the equation
$\frac{1}{X(1-q^{-1}X)}s-Y_{2}s^{2}-\frac{s}{X(1-q^{-1}X-\frac{1}{q^{-1}X(1-q^{-2}X)}s+...)}=0,$
We rewrite the last term as
$\frac{s}{X(1-q^{-1}X-\frac{1}{q^{-1}X(1-q^{-2}X)}s+...)}$
$=$ $\frac{s}{X(1-q^{-1}X)}\cdot\frac{1}{1-\frac{1}{(1-q^{-1}X)q^{-1}X(1-q^{-2}X)}s+...)}$
$=$ $\frac{s}{X(1-q^{-1}X)}\cdot\left(1+\frac{1}{(1-q^{-1}X)q^{-1}X(1-q^{-2}X)}s+...\right)$
The first term cancels the first term in the equation above. To cancel the second term we need to have
$Y_{2}=-\frac{1}{X^{2}q^{-1}(1-q^{-1}X)^{2}(1-q^{-2}X)}.$
It is clear that the period equation (2.2) can be written in terms of the coefficients $Y_{n}(X).$ First, some notation. For a power series $B=\sum B_{n}s^{n},$ recall we write $[s^{n}]B$ for the coefficient $B_{n}$, and let us define the truncation $\hat{B}_{\leq n}:=\sum_{k\leq n}B_{k}s^{k}.$
Then we can write the difference equation as the recursion $Y_{0}=1-X$ and
$Y_{n+1}=[s^{n}]\frac{-1}{XY_{\leq n}(q^{-1}X)}$
for $n\geq 0.$
Note that $Y(X)\in 1+X\mathbb{C}[X]+s\mathbb{C}[q^{\pm 1},X^{-1}]\llbracket X\rrbracket\llbracket s\rrbracket$, hence $\log(Y(X))\in X\mathbb{C}\llbracket X\rrbracket+s\mathbb{C}[q^{\pm 1},X^{-1}]\llbracket X\rrbracket\llbracket s\rrbracket$. In particular, $\log(Y(X))$ has a well-defined residue and constant term with respect to $X$. Let’s compute:
$Y=1-X+Y_{1}s+Y_{2}s^{2}+...=(1-X)\left(1+\frac{Y_{1}}{1-X}s+\frac{Y_{2}}{(1-X)}s^{2}+...\right),$
so
$\log(Y)=\log(1-X)+\log\left(1+\frac{Y_{1}}{1-X}s+\frac{Y_{2}(1-X)-Y_{1}^{2}/2}{(1-X)^{2}}s^{2}+...\right).$
Next write
$\underline{a}(z,q)={\sf Const}(\log Y(X))=\sum{R}_{n}(q)s^{n}=\sum{R}_{n}(q)q^{\frac{n}{2}}z^{n}=:\sum\underline{a}_{n}(q)z^{n},$
where we have recalled $s=q^{\frac{1}{2}}z$ and defined the coefficients $\underline{a}_{n}(q)$ of $\underline{a}(z,q).$
Then $\underline{a}_{0}(q)=0.$ Next $Y_{1}=-\frac{1}{X(1-q^{-1}X)}$ gives $R_{1}=-q^{-1}-1,$ and $\underline{a}_{1}(q)=-q^{\frac{1}{2}}-q^{-\frac{1}{2}},$ as in Example 2.15. From the above form of $Y_{2},$ we also find
$\underline{a}_{2}(q)=-q^{2}-\frac{7}{2}q-6-\frac{7}{2}q^{-1}-q^{-2}.$
Using this method, it is simple to find the quantum period to any desired order. To $O(z^{3})$ we find
$\begin{array}[]{l}\underline{a}(z,q)=\left({-q^{\frac{1}{2}}-q^{-\frac{1}{2}}}\right)z+\left(-q^{2}-\frac{7}{2}q-6-\frac{7}{2}q^{-1}-q^{-2}\right)z^{2}+\\[5.0pt] \ \left(-q^{\frac{9}{2}}-3q^{\frac{7}{2}}-12q^{\frac{5}{2}}-\frac{88}{3}q^{\frac{3}{2}}-48q^{\frac{1}{2}}-48q^{-\frac{1}{2}}-\frac{88}{3}q^{-\frac{3}{2}}-12q^{-\frac{5}{2}}-3q^{-\frac{7}{2}}-{q^{-\frac{9}{2}}}\right)z^{3}+\ldots\end{array}$
As remarked in Example 1.2, this $q$-refines the classical period. Moreover, it agrees with $a(z,q)$ in Example 2.15. By Theorem 2.23 below, this is not a coincidence.
2.7. Equality of definitions for the quantum A-period
Recall from §2.2 that $p(X,Y)=1-f(X,Y)$ gets quantized to an operator which acts on a function $\Phi(X)$ by $X\cdot\Phi(X)=X\Phi(X)$ and $Y\cdot\Phi(X)=\Phi(qX)$, and the wavefunction $\Psi(X)$ is defined by $p\cdot\Psi=0$. We have defined
$a(z,q)={\sf Const}_{X,Y}(\log p(X,Y))=-\sum_{n>0}\frac{1}{n}{\sf Const}_{X,Y}(f^{n})$
and by Corollary 2.22,
$\underline{a}(z,q)={\sf Const}_{X}(\log Y(X)),\quad Y(X)=\frac{\Psi(qX)}{\Psi(X)},$
so that the action of $Y$ on $\Psi(X)$ is given by multiplication with $Y(X)$. Note that for simplicity we have hidden the dependencies on $t$, $z$ and $q$ here.
Theorem 2.23.
We have $a(z,q)=\underline{a}(z,q)$.
Proof.
Substituting $X=xt$ and $Y=yt$, we get $f(X,Y,z)=tf(x,y,z)|_{z^{\beta}=t^{\beta\cdot E}z^{\beta}}$. E.g., for $\mathbb{P}^{2}$ we have $f=X+Y+zX^{-1}Y^{-1}=t(x+y+t^{3}zx^{-1}y^{-1})$. Applying $D$ as in §1.2, the claim is equivalent to
$D{\sf Const}_{X,Y}(\log p(X,Y))=1-{\sf Const}_{x,y}\left(\frac{1}{1-tf(x,y)}\right)=D{\sf Const}_{X}(\log Y).$
Note that $g(t,y):=(1-tf)^{-1}\in(R[y]\llbracket y^{-1}\rrbracket)\llbracket t\rrbracket$ for $R=\mathbb{C}[\textup{NE}(X)][x]\llbracket x^{-1}\rrbracket$. By Convention 2.5, we can assume that $f(X,Y)$ does not contain any terms of $Y$-order $>1$. In other words, $\operatorname{ord}_{Y}(f)=1$, with Definition A.2. By Lemma A.3, $y(x,f)$ is a Laurent series of order $1$ in $f$. Then we can use Lemma A.6 to get
${\sf Const}_{x,y}\left(\frac{1}{1-tf(x,y)}\right)$ $=$ ${\sf Const}_{x,f}\left(\frac{1}{1-tf}\frac{f\partial_{f}y(x,f)}{y(x,f)}\right)$
$=$ ${\sf Const}_{x,f}\left(\left(1+tf+t^{2}f^{2}+\ldots\right)\frac{f\partial_{f}y(x,f)}{y(x,f)}\right)$
$=$ $\sum_{k=0}^{\infty}{\sf Const}_{x}\left([f^{-k}]\frac{f\partial_{f}y(x,f)}{y(x,f)}\right)t^{k}.$
Here $[f^{-k}]$ is the operator that takes the coefficient of $f^{-k}$ in the Laurent series right of it. In the last equality we have used that $y(x,f)$ does not depend on $t$. Now recall $p=1-tf(x,y)=0$. Hence we have $f=1/t$ and $\partial_{f}y(x,f)=-t^{2}\partial_{t}y(x,t)$. Inserting this into the equation above gives
${\sf Const}_{x,y}\left(\frac{1}{1-tf(x,y)}\right)$ $=$ $\sum_{k=0}^{\infty}{\sf Const}_{x}\left([t^{k}]\frac{f\partial_{f}y(x,f)}{y(x,f)}\right)t^{k}$
$=$ ${\sf Const}_{x}\left(\frac{f\partial_{f}y(x,f)}{y(x,f)}\right)$
$=$ ${\sf Const}_{x}\left(-\frac{t\partial_{t}y(x,t)}{y(x,t)}\right)$
$=$ $-{\sf Const}_{x}\left(t\tfrac{d}{dt}\log y\right),$
Note that taking the logarithmic derivative $t\tfrac{d}{dt}$ has the same effect as $D$, since $t$ captures the total $z$-degree. Using $\log Y=\log y+\log t$ we get
$D{\sf Const}_{X}(\log Y)=1+{\sf Const}_{x}(D\log y)=1-{\sf Const}_{x,y}\left(\frac{1}{1-tf(x,y)}\right),$
as claimed. ∎
2.8. Natural $q$-refinements of mirror maps
We recall that in [GRZ], the open mirror map was defined by $U=xe^{a(z)},$ relating complex and symplectic open moduli $x$ and $U$ via $a(z)$, the A-period of the mirror curve, which depended on the closed complex modulus, $z$. Moreover $z$ and the symplectic modulus $Q$ were related by $Q=ze^{-a(z)},$ and the function $M(Q)=e^{-a(z)}$ was given enumerative interpretation. Here we define some straightforward $q$-analogues.
Lemma 2.24.
Let $A$ be a Noetherian algebra over $A_{0}$ with an ideal $I$ so that the composition $A_{0}\to A\to A/I$ is an isomorphism. The formal completion of $A$ in the ideal $I$ is denoted $\widehat{A}$. Suppose $\phi\colon\widehat{A}\to\widehat{A}$ is an $A_{0}$-algebra homomorphism satisfying $\phi(I)\subseteq I$ and inducing an isomorphism of $\phi\colon\widehat{A}/I^{2}\to\widehat{A}/I^{2}$ then $\phi$ is an isomorphism.
Proof.
By assumption, $\phi$ respects the filtration by powers of $I$. We first show by induction over $k$ that $\phi$ induces an isomorphism on $I^{k}/I^{k+1}$. The base case $I^{0}/I=A_{0}$ is given. The multiplication map $m\colon I/I^{2}\otimes I^{n}/I^{n+1}\to I^{n+1}/I^{n+2}$ is easily seen to be surjective. The induction hypothesis gives an isomorphism on the source of $m$ which implies it on the target, inducing the induction step.
By another induction over $k$, we now show that $\phi$ induces an isomorphism on $A/I^{k}$ which is straightforward when applying the five lemma to the $\phi$-induced endomorphism of the exact sequence $0\to I^{k-1}/I^{k}\to A/I^{k}\to A/I^{k-1}\to 0$.
[…]
Proposition 2.25.
Mapping monomials by $\phi\colon Q^{\beta}\mapsto z^{\beta}e^{-(\beta\cdot E)a(-z,q)}$ defines an automorphism of $\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket\textup{NE}(\mathbb{P})\rrbracket.$
Proof.
It is straightforward to check that $\phi(Q^{\beta})\cdot\phi(Q^{\beta^{\prime}})=\phi(Q^{\beta+\beta^{\prime}})$, so $\phi$ gives a well-defined endomorphism. If $I$ denotes the ideal generated by all monomials $Q^{\beta}$ in the ring $A=\mathbb{C}[q^{\pm\frac{1}{2}}][\textup{NE}(\mathbb{P})]$ then $A$, $I$, $\widehat{A}=\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket\textup{NE}(\mathbb{P})\rrbracket$ and $\phi$ satisfy the assumptions of Lemma 2.24 which gives the assertion. ∎
Definition 2.26.
Put $U=xe^{a(-z,q)},$ a relation between two formal parameters $x$ and $U$ via the quantum period of Definition 2.14. We define $Q^{\beta}(z,q)=z^{\beta}e^{-(\beta\cdot E)a(-z,q)}$ and by Proposition 2.25 obtain the inverse expressions $z^{\beta}(Q,q)$. We define the function
(2.3) $M(Q,q):=e^{-a(-z(Q,q),q)}=\frac{U}{x}=\left(\frac{z^{\beta}(Q,q)}{Q^{\beta}}\right)^{-1/(\beta\cdot E)}$
where the last equality holds for every non-zero class $\beta\in\operatorname{NE}(\mathbb{P})$. The series $M(Q,q)$ will be given an enumerative interpretation in Theorem 4.2.
Example 2.27.
In the case of $\mathbb{P}^{2}$, we have
$\begin{array}[]{l}M(Q,q)=1-\left(q^{\frac{1}{2}}+q^{-\frac{1}{2}}\right)Q+\left(q^{2}+q+1+q^{-1}+q^{-2}\right)Q^{2}\\[5.0pt] \ \ -\left(q^{\frac{9}{2}}+3q^{\frac{7}{2}}+4q^{\frac{5}{2}}+4q^{\frac{3}{2}}+4q^{\frac{1}{2}}+4q^{-\frac{1}{2}}+4q^{-\frac{3}{2}}+4q^{-\frac{5}{2}}+3q^{-\frac{7}{2}}+q^{-\frac{9}{2}}\right)Q^{3}+\cdots,\end{array}$
as computed in Example 1.4.
3. Quantum periods, broken lines and descendant invariants
In this section, we give a combinatorial construction of the quantum period (Definition 2.14) via broken lines in a consistent scattering diagram. In §4 we will use this description to prove that the primitive theta function, defined using broken lines, equals the open mirror map, also after $q$-refinement. Moreover, we discuss a relationship with descendant Gromov-Witten invariants with $\lambda$-class insertions.
3.1. Toric degenerations and scattering diagrams
In this subsection and the next we summarize the definition of scattering diagrams and broken lines with closed moduli $z^{\beta}\in\mathbb{C}[\textup{NE}(\mathbb{P})]$ and non-commutative variables $X,Y$ ($q$-refined). For more details see [Bou][CPS][GRZ].
Let $\mathbb{P}$ be a smooth projective Fano surface and let ${E}$ be a smooth anticanonical divisor, so that $(\mathbb{P},{E})$ is a smooth log Calabi-Yau pair. As in §2.1, if $\mathbb{P}$ is not toric, consider a $\mathbb{Q}$-Gorenstein degeneration of $\mathbb{P}$ to a Gorenstein toric Fano variety $\mathbb{P}_{\Sigma^{\prime}}$ or a smooth toric semi-Fano variety $\mathbb{P}_{\Sigma}$. Let $\Delta^{\star}$ be the reflexive polygon given by the convex hull of the ray generators of $\Sigma^{\prime}$ (or $\Sigma$) and let $\Delta$ be the polar dual polytope. Then $\Delta$ is the toric polytope of $\mathbb{P}_{\Sigma^{\prime}}$. The lattice points of $\Delta$ define an embedding of the projective toric variety $\mathbb{P}_{\Sigma^{\prime}}$ into $\mathbb{P}^{|\Delta\cap\mathbb{Z}^{2}|-1}$. The polytope $\Delta$ is reflexive. In particular it contains a unique interior lattice point.
Consider the central subdivision of $\Delta$ by adding, for each vertex, an edge that connects the vertex with the interior point. We can smooth the boundary of $\Delta$ by adding affine singularities on these new edges. This induces a toric degeneration of $(\mathbb{P},{E})$, i.e., a degeneration into a union of toric pieces such that the family is strictly semi-stable away from a codimension $2$ subset. The irreducible components of the central fiber of the toric degeneration correspond to the maximal cells of the central subdivision of $\Delta$. In other words, the central subdivision of $\Delta$ is an affine manifold with singularities that is the intersection complex of the toric degeneration.
By dualizing we get the dual intersection complex $B$. This is an affine manifold with singularities that has a natural polyhedral decomposition (dual to the central subdivision of $\Delta$). It has a unique bounded maximal cell (dual to the interior lattice point of $\Delta$) that is isomorphic to $\Delta^{\star}$. The other maximal cells are unbounded. The affine singularities are such that all unbounded edges of $B$ are parallel, so that there is a unique unbounded direction $m_{\text{out}}$. We write $\textup{Sing}(B)$ for the singular locus of $B$, which is a finite set of points, and $B_{0}:=B\setminus\textup{Sing}(B)$.
Example 3.1.
For $(\mathbb{P}^{2},E)$, with $E\subset\mathbb{P}^{2}$ an elliptic curve, a toric degeneration into a union of three $\mathbb{P}(1,1,3)$’s is given by
$\mathcal{X}$ $=$ $\{XYZ=t^{3}(W+X^{3}+Y^{3}+Z^{3})\}\subset\mathbb{P}(1,1,1,3)\times\mathbb{A}_{t}^{1}\rightarrow\mathbb{A}_{t}^{1}$
$\mathcal{D}$ $=$ $\{W=0\}\subset\mathcal{X}$
Figure 3.1 shows the corresponding intersection complex (= subdivision of the moment polytope) and the dual intersection complex (sometimes called tropicalization).
$\times$$\times$$\times$
$\times$$\times$$\times$
Figure 3.1. The intersection complex (left) and dual intersection complex (right) of $(\mathbb{P}^{2},E)$.
A multi-valued piecewise linear function on an open subset $U\subset B_{0}=B\setminus\textup{Sing}(B)$ is given by a collection of piecewise linear functions on an open covering $\{U_{i}\}$ of $U$ which differ only by affine linear functions on overlaps. Choose a cyclic labeling $\rho_{1},\ldots,\rho_{r+2}$ of the rays of $\Sigma$, hence of the boundary lattice points $m_{1},\ldots,m_{r+2}$ of $\Delta^{\star}$. For each $k=1,\ldots,r$ define a multi-valued piecewise linear function $\varphi_{k}$ on $B_{0}$ as follows. For each maximal cell $\sigma$ of the polyhedral decomposition of $B$ (there are $r+3$ such cells, $1$ bounded and $r+2$ unbounded), let $V_{\sigma}$ be the interior of $\sigma$. For each $i=1,\ldots,r+2$, let $U_{i}\subset B_{0}$ be a neighborhood of $m_{i}$ such that its closure $\overline{U}_{i}$ contains both affine singularities adjacent to $m_{i}$. Then $\{U_{i}\}\cup\{V_{\sigma}\}$ gives an open covering of $B_{0}$. Define $\varphi_{k}|_{V_{\sigma}}=0$ for all $\sigma$ and $\varphi_{k}|_{U_{i}}=0$ for all $i\neq k$. Define $\varphi_{k}|_{U_{k}}$ by the conditions that it is 00 on $\Delta^{\star}$ and has slope $1$ along the ray $\rho_{k}$. Note that the bend locus of $\varphi_{k}$ is given by $\rho_{k}$ and the two line segments between $m_{k}$ and its adjacent affine singularities.
The dual intersection complex $B$ defines a scattering diagram (see [Bou]) with initial rays emanating from the affine singularities. The rays in the scattering diagram carry functions in $\mathbb{C}[\textup{NE}(\mathbb{P})]\llbracket x^{m}\rrbracket$, where $m$ is the tangent direction of the ray. The scattering algorithm gives a unique consistent scattering diagram.
For $q$-refined scattering (see [Bou]), the commutative variables $x,y$ are replaced by non-commutative variables $X,Y$ with commutation law $XY=q^{-1}YX$.
Example 3.2.
For $(\mathbb{P}^{2},E)$, the multi-valued piecewise linear function $\varphi$ and the scattering diagram are shown in Figure 3.2.
$\times$$\times$$\times$$\ \ \ \ \varphi=0$$\varphi=x-2y-1$$\varphi=-2x+y-1$$\varphi=0$$\varphi=0$$\varphi=0$$\varphi=0$$\varphi=0$$\varphi=0$
$\times$$\times$$\times$
Figure 3.2. The multi-valued piecewise linear function (left) and the scattering diagram (right) for $(\mathbb{P}^{2},E)$.
3.2. Theta functions and quantum periods
A broken line is a piecewise affine linear embedding $\mathfrak{b}:(-\infty,0]\rightarrow B_{0}=B\setminus\textup{Sing}(B)$ with $\mathfrak{b}(0)=P$. The domains of affine linearity carry functions that are Laurent monomials in a local coordinate $x^{m}$ with coefficients in $\mathbb{C}[\textup{NE}(\mathbb{P})]$, where $m$ is the tangent direction of (the graph of) $\mathfrak{b}$. A broken line can “break” (i.e., change its domain of affine linearity) only if it intersects a ray of the scattering diagram. In this case, the monomials of the adjacent domains of affine linearity are related by a wall crossing morphism.
When a broken line $\mathfrak{b}$ crosses the bend locus of $\varphi_{k}$ such that $\varphi_{k}$ changes slope along $\mathfrak{b}$ by $\kappa\in\mathbb{Z}$, then $\mathfrak{b}$ picks up a factor of $z^{\kappa\beta_{k}}$. Moreover, the initial rays starting in an affine singularity contained in $\overline{U}_{k}$ but emanating in the opposite direction carry a factor of $z^{\beta_{k}}$. If a broken line $\mathfrak{b}$ breaks at such a ray or at any ray that is produced from it later in the scattering procedure, then $\mathfrak{b}$ picks up a factor of $z^{\beta_{k}}$.
The consistent scattering diagram supports theta functions $(\vartheta_{1})_{P}$ defined as a sum over broken lines with primitive incoming monomial $x^{m_{\text{out}}}$ and ending in a prescribed point $P\in B_{0}$. Note that the incoming monomial has $z$-degree zero. Hence, it can be seen as a tangent direction lying on the graph of $\varphi_{1}\times\cdots\times\varphi_{k}$. The theta functions $(\vartheta_{1})_{P}$ only depend on the chamber of the scattering diagram in which $P$ lies, and are independent of the point within a chamber. We write $(\vartheta_{1})_{0}$ for the theta function inside the bounded maximal cell of $B$, which is a chamber of the scattering diagram by convexity of $\Delta^{\star}$. We write $(\vartheta_{1})_{\infty}$ for the theta function in an unbounded chamber of the consistent scattering diagram. In this case, the endpoint $P$ will be “infinitely” far away (we take a formal limit here) from $\Delta^{\star}$, hence the notation.
For $q$-refined theta functions, the commutative variables $x,y$ are replaced by non-commutative variables $X,Y$ with commutation law $XY=q^{-1}YX$.
Example 3.3.
For $(\mathbb{P}^{2},E)$, Figure 3.3 shows the broken lines that give
$\vartheta_{1}(X,Y,z)_{0}=X+Y+zq^{\frac{1}{2}}X^{-1}Y^{-1}.$
$\times$$\times$$\times$
$\times$$\times$$\times$
Figure 3.3. Broken lines in the scattering diagram defining $\vartheta_{1}$ in the central chamber (left) and at infinity up to order $2$ (right).
By [Man], Proposition 2.15, the structure constants of $q$-refined theta multiplication do not depend on the chamber. In particular ${\sf Const}(\vartheta_{1}(q)^{n})$ does not depend on the chamber. Hence, the following definition makes sense.
Definition 3.4.
The $q$-refined classical period of $\vartheta_{1}(q)$ is
$\pi(z,q)=\sum_{n>0}{\sf Const}_{X,Y}(\vartheta_{1}(q)^{n}),$
independent of the chamber in which $\vartheta_{1}(q)$ is defined.
Example 3.5.
For $(\mathbb{P}^{2},E)$ we have
$\pi(z,q)=(3q^{\frac{1}{2}}+3q^{-\frac{1}{2}})z+(6q^{2}+21q+36+21q^{-1}+6q^{-2})z^{2}+\ldots.$
Lemma 3.6.
In the central chamber of the scattering diagram, we have
$\vartheta_{1}(X,Y,z)_{0}=f(X,Y,z),$
where $f(X,Y,z)$ is the Laurent polynomial from Definition 2.1, with variables $x,y$ replaced by non-commutative variables $X,Y$ with commutation law $XY=q^{-1}YX$.
Proof.
Choose the endpoint $P$ close to the origin. Then for each boundary lattice point $p_{k}$ of $\Delta^{\star}$ we get broken line with image $P+\mathbb{R}_{\geq 0}p_{k}$ that does not break at any ray. In other words, this broken line is actually a straight line. If $p_{k}$ is lies in the interior of an edge of $\Delta^{\star}$, then we have more broken lines with image $P+\mathbb{R}_{\geq 0}p_{k}$, obtained as follows. Let $P_{k}$ be the point where $P+\mathbb{R}_{\geq 0}p_{k}$ intersects the boundary of $\Delta^{\star}$. Let $m_{k}$ and $n_{k}$ be the lattice distances of $p_{k}$ from the two vertices of the edge. Choose $0<l\leq\min(m,n)$ and choose $l$ distinct affine singularities on each side of $P_{i}$. Then we can break at the rays emanating from these singularities. This breaking all happens simultaneously at the same point $P_{k}\in B$. Since we have chosen the same number $l$ on both sides, the breaking does not change the image of the broken line. The choice of affine singularities corresponds to the choice of $0<i_{1}<\ldots<i_{l}<m_{k}$ and $0<j_{1}<\ldots<j_{l}<n_{k}$ in the definition of $f$ (Definition 2.1). The $z$-dependence in the definition of $f$ is exactly the one we get from breaking at these rays. Hence $(\vartheta_{1})_{P}$ close to the origin agrees with $f$. By [Man], Theorem 2.14, $(\vartheta_{1})_{P}$ is constant in a chamber. Hence we have $(\vartheta_{1})_{0}=f$ in the whole central chamber. ∎
Recall from §1.2 that we defined an operation $D$ on $\mathbb{C}[\textup{NE}(\mathbb{P})]$ by $Dz^{\beta}=(\beta\cdot D)z^{\beta}$.
Proposition 3.7.
We have
$\pi(z,q)=-Da(z,q),$
Proof.
By Definition 2.14, we have
$-a(z,q)=\sum_{n>0}\frac{1}{n}{\sf Const}_{X,Y}(f^{n}).$
By the definition of $f$, every monomial in ${\sf Const}_{X,Y}(f^{n})$ takes the form $c(q)z^{\beta}$ with $\beta\cdot E=n$. By the definition of $D$ we have $Dz^{\beta}=(\beta\cdot E)z^{\beta}$. Hence, $D{\sf Const}_{X,Y}(f^{n})=n{\sf Const}_{X,Y}(f^{n})$. Hence, applying $D$ to the equation above gives
$-Da(z,q)=\sum_{n>0}{\sf Const}_{X,Y}(f^{n}),$
which by Definition 3.4 and Lemma 3.6 agrees with $\pi(z,q)$. ∎
It was shown in [GRZ], Proposition 3.5, that the broken line expansion of the theta function at infinity, denoted $(\vartheta_{1})_{\infty}$, is a Laurent series in a single variable $y=x^{m_{\text{out}}}$ corresponding to the unique asymptotic direction $m_{\text{out}}$. Consequently, the monomials in $(\vartheta_{1})_{\infty}$ are commutative. The series $(\vartheta_{1})_{\infty}$ has the following enumerative interpretation.
Proposition 3.8.
Let $q=e^{i\hbar}$. Then
$\vartheta_{1}(y,z,q)_{\infty}=y+\sum_{\beta\in\textup{NE}(\mathbb{P})}\sum_{g\geq 0}\frac{1}{\beta\cdot E-1}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}y^{-(\beta\cdot E-1)},$
where $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ is the $2$-marked logarithmic Gromov-Witten invariant of genus $g$ with $\lambda_{g}$-insertion. Moreover, $\vartheta_{1}(y,z,q)_{\infty}$ is a formal Laurent series in $y$ over the ring $\mathbb{C}[q^{\pm\frac{1}{2}}][\textup{NE}(\mathbb{P})]$.
Proof.
By [GRZ], Theorem 4.14, we have
$\vartheta_{1}(y,z,q)_{\infty}=y+\sum_{\beta\in\textup{NE}(\mathbb{P})}\sum_{g\geq 0}R^{\textup{trop}}_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}y^{-(\beta\cdot E-1)},$
where $R^{\textup{trop}}_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)$ is the tropical analogue of $R^{\textup{trop}}_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)$, defined as a count (with $q$-refined multiplicity) of rational tropical curves on the dual intersection complex $B$ of $(\mathbb{P},E)$ with an unbounded leg of weight $\beta\cdot E-1$ through a prescribed point, an unbounded leg of weight $1$ and no other unbounded legs. By [Gra2], Theorem 6.2, we have
$R^{\textup{trop}}_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)=(\beta\cdot E-1)R_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta),$
and by [GRZ], Theorem 5.4 we have
$(\beta\cdot E-1)R_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)=\frac{1}{\beta\cdot E-1}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta).$
The three displayed equations together yield the claimed equality. The $y^{-k}$-coefficient of $\vartheta_{1}(y,z,q)_{\infty}$ is a sum over finitely many $\beta\in\textup{NE}(\mathbb{P})$, those with $\beta\cdot E=k+1$. For each $\beta\in\textup{NE}(\mathbb{P})$, the coefficient $R^{\textup{trop}}_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)$ is a Laurent polynomial in $q^{\pm\frac{1}{2}}$ by the definition of $q$-refined tropical multiplicity (see [GRZ], Definition 4.6). Hence, $\vartheta_{1}(y,z,q)_{\infty}$ is a formal Laurent series over the ring $\mathbb{C}[q^{\pm\frac{1}{2}}][\textup{NE}(\mathbb{P})]$. ∎
Example 3.9.
For $(\mathbb{P}^{2},E)$, Figure 3.3 shows the broken lines that give
$\begin{array}[]{l}\vartheta_{1}(y,z,q)_{\infty}=y+\left(q^{\frac{1}{2}}+q^{-\frac{1}{2}}\right)zy^{-2}+\left(q^{2}+q+1+q^{-1}+q^{-2}\right)z^{2}y^{-5}\\[5.0pt] \ \ +\left(q^{\frac{9}{2}}+3q^{\frac{7}{2}}+4q^{\frac{5}{2}}+4q^{\frac{3}{2}}+4q^{\frac{1}{2}}+4q^{-\frac{1}{2}}+4q^{-\frac{3}{2}}+4q^{-\frac{5}{2}}+3q^{-\frac{7}{2}}+q^{-\frac{9}{2}}\right)z^{3}y^{-8}+\cdots.\end{array}$
3.3. Quantum descendant invariants
The mirror symmetry proposal of [CCGGK] states that a Laurent polynomial $f$ is mirror dual to $\mathbb{P}$ if the classical period of $f$ agrees with the regularized GW period [^4] This term was referred to as a quantum period in [CCGGK], a terminology we need to avoid here. of $\mathbb{P}$, which is defined as a generating function of descendant invariants. We give a $q$-refinement of this statement and prove it using tropical geometry.
Definition 3.10.
Let $q=e^{i\hbar}$. The $q$-refined regularized GW period of $\mathbb{P}$ is
$\widehat{G}(z,q)=1+\sum_{\beta\in\textup{NE}(\mathbb{P})}\sum_{g\geq 0}(\beta\cdot E)!D_{g,1}(\mathbb{P},\beta)\hbar^{2g}z^{\beta},$
with descendant Gromov-Witten invariants
$D_{g,1}(\mathbb{P},\beta)=\int_{\mathcal{M}_{g,1}(\mathbb{P},\beta)}(-1)^{g}\lambda_{g}\psi^{\beta\cdot{E}-2}\textup{ev}^{\star}[\textup{pt}],$
where $\lambda_{g}=c_{g}(\mathbb{E})$ are the Chern classes of the Hodge bundle.
Remark 3.11.
This table summarizes our notation.
B-side A-side
classical period GW period
quantum period $q$-refined GW period
Proposition 3.12.
We have
$\pi(z,q)=\widehat{G}(z,q).$
Proof.
By the multiplication rule of theta functions, ${\sf Const}(\vartheta_{1}(q)^{m})$ is a sum over $m$-tuples of broken lines that are balanced in the endpoint. In the central chamber the broken lines are simply straight lines with ending monomials $z^{m_{\rho}}$. So ${\sf Const}(\vartheta_{1}(q)^{m})$ is the count of balanced $m$-tuples of broken lines. But such an $m$-tuple is nothing but a tropical curve with a fixed point lying on a vertex $V$ with valency $m$. The tropical curve has automorphism group of order $m!$ and is counted with a $q$-refined multiplicity defined in [BlSh], see [KSU1], §1.2. By [KSU1], Theorem A, with $n=1$ and $k_{1}=|\Delta^{\circ}|-2$, the tropical count equals the descendant invariants $D_{g,1}(\mathbb{P},\beta)$, such that
${\sf Const}(\vartheta_{1}(q)^{m})=\sum_{\beta:\beta\cdot{E}=m}\sum_{g\geq 0}m!D_{g,1}(\mathbb{P},\beta)\hbar^{2g}z^{\beta}.$
Remark 3.13.
Proposition 3.12 is proved in genus zero in [Joh] and for non-smooth divisors (Looijenga pairs) in [KSU2]. Propositions 3.7 and 3.12 together yield a relation between $2$-marked invariants $R^{\textup{trop}}_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)$ and descendant invariants $D_{g,1}(\mathbb{P},\beta)$. In genus zero, this relation was also deduced in [Joh].
Theorem 3.14.
We have
$-Da(z,q)=\widehat{G}(z,q).$
Proof.
The statement follows from Propositions 3.7 and 3.12. ∎
Remark 3.15.
The non-$q$-refined version of Theorem 3.14 was proved in [Giv] using
$\widehat{G}(z,q=1)=\sum_{\beta\in\textup{NE}(\mathbb{P})}\frac{(-K_{\mathbb{P}}\cdot\beta)!}{(D_{1}\cdot\beta)!\cdots(D_{r}\cdot\beta)!}z^{\beta}.$
4. The $q$-refined open mirror map is the quantum theta function
We prove the identification of the quantum theta function asymptotic broken line expansion and the q-refined open mirror map in this section. As a consequence, we obtain the enumerative expression for the quantum period.
Proposition 4.1.
Let $A=\bigoplus_{k\geq 0}A_{k}$ be a finitely generated graded algebra over $A_{0}$ and let $\widehat{A}$ denote the formal completion of $A$ in the ideal $\bigoplus_{k\geq 1}A_{k}$. Let $\theta(y)\in A\llbracket y^{-1}\rrbracket[y]$ be a formal Laurent series of the form $\theta(y)=y+\sum_{k\geq 0}a_{k}y^{-k}$ with $a_{k}\in A_{k+1}$. The series $\widetilde{y}=-\exp\left({-\sum_{k>0}\frac{(-1)^{k}}{k}[y^{0}]\theta^{k}}\right)$ gives a well-defined element of $\widehat{A}$ that satisfies
$\theta(\tilde{y})=-1.$
Moreover, $\tilde{y}$ uniquely determines the coefficients $a_{k}$, hence it determines a function $\theta(y)$ of the given form.
Proof.
Define $f(y)=\theta(y^{-1})$ and let $A_{r}$ be the degree $r$ graded piece of $A$. Then $f(y)$ is a formal Laurent series in $y$ of the form $f(y)=y^{-1}+\sum_{k\geq 0}f_{k}y^{k}$ such that $f_{k}\in A_{k+1}$. Recall from §1.2 that we write $[y^{k}]f$ for the coefficient of $y^{k}$ in $f$. The power series $h(y):=yf(y)=1+\sum_{k>0}f_{k-1}y^{k}\in A\llbracket y\rrbracket$ has the convenient property $[y^{k}]h^{l}\in A_{k}$ for every $l\geq 0$. In particular, $[y^{k}](yf)^{k}\in A_{k}$, so $[y^{0}]f^{k}\in A_{k}$ and
(4.1) $[y^{-1}]f^{k}\in A_{k-1}.$
We apply Proposition A.12 to $h$ to have
$\exp\left(\sum_{k>0}\frac{1}{k}[y^{k}]h^{k}w^{k}\right)=\sum_{k>0}\frac{1}{k}[y^{k-1}]h^{k}w^{k-1},$
with a formal variable $w$. We multiply both sides with $w$, insert $h=yf$ and define
$g(w):=w\exp\left(\sum_{k>0}\frac{1}{k}[y^{0}]f^{k}w^{k}\right)=\sum_{k>0}\frac{1}{k}[y^{-1}]f^{k}w^{k}.$
Comparing the very last expression with Lemma A.1 shows [^5] While Lemma A.1 is only stated for complex coefficient, we can embed our coefficient ring $A\llbracket y\rrbracket[y^{-1}]$ into $\mathbb{C}$ before using the lemma. that $g(w^{-1})$ is the compositional inverse of $f(y)$. That is, we have $f(g(w^{-1}))=w$, or equivalently $f(g(w))=w^{-1}$.
We write $g(w)=\sum_{k>0}g_{k}w^{k}$ and have $g_{k}=[w^{k}]g=\frac{1}{k}[y^{-1}]f^{k}$. By (4.1), $g_{k}\in A_{k-1}$. This allows us to specialize $w$ to any fixed value $c$ and the resulting series $g(c)$ gives a well-defined element in the formal completion $\widehat{A}$.
We set $w=-1$. Then $f(g(w))=w^{-1}$ becomes $f(g(-1))=-1$. It is the substitution of an element of $\widehat{A}$ for $y$ in an element of $A\llbracket y\rrbracket[y^{-1}]$ and reads
(4.2) $f\left(-\exp\left({\sum_{k>0}\frac{(-1)^{k}}{k}[y^{0}]f^{k}}\right)\right)=-1.$
We recognize that $\widetilde{y}=\frac{1}{g(-1)}$ and the assertion follows from (4.2) by inserting the definition of $f$ that gives $\theta(y)=f(y^{-1})$.
By Lemma A.1, the compositional inverse $g(w)$ uniquely determines $\theta(y)$. We have argued above that $g_{k}\in A_{k-1}$. Hence, the $w$-exponent in $g$ only captures the degree of $g_{k}$ in the graded ring $A$. In turn, we can recover $g(w)$ from any fixed value of $w$. In particular, we can recover $g(w)$ from $\tilde{y}=\frac{1}{g(-1)}$ and this determines $\theta(y)$ under the requirement that $\theta(y)$ is of the given form. ∎
Using the quantum A-period $a(z,q)$ from Definition 2.14, in Definition 2.26 we defined a natural $q$-refinement of the open and closed mirror map by
(4.3) $M(Q,q)=e^{-a(-z(Q),q)},\quad Q^{\beta}(z,q)=z^{\beta}e^{-(\beta\cdot{E})a(-z,q)}.$
Theorem 4.2.
The insertion of $Q^{\beta}(z,y)=z^{\beta}\cdot(-1/y)^{\beta\cdot{E}}$ into $M(Q,q)$ gives the identity
$\vartheta_{1}(y,z,q)_{\infty}=yM(Q,q).$
Proof.
The ring $A:=\mathbb{C}[q^{\pm\frac{1}{2}}][\textup{NE}(\mathbb{P})]$ is graded by $\deg(z^{\beta})=\beta\cdot E$. By Proposition 3.8, $\vartheta_{1}(y,z,q)_{\infty}$ is a formal Laurent series over $A$ of the form $\vartheta_{1}(y,z,q)_{\infty}=y+\sum_{k\geq 0}a_{k}y^{-k}$ with $a_{k}\in A_{k+1}$. Then Proposition 4.1 shows that the insertion of
$y(z,q):=-\exp\left({-\sum_{k>0}\frac{(-1)^{k}}{k}[y^{0}]\vartheta_{1}^{k}}\right)\in\mathbb{C}[q^{\pm\frac{1}{2}}]\llbracket\textup{NE}(\mathbb{P})\rrbracket$
gives $\vartheta_{1}(y(z,q),z,q)_{\infty}=-1$. Comparing $y(z,q)$ with Proposition 3.7 shows
$y(z,q)=-\exp({a(-z,q)}).$
Comparing with $M(Q,q)=\exp\big{(}{-a(-z(Q,q),q)}\big{)}$ from (4.3), yields
(4.4) $y(z,q)=\frac{-1}{M(Q(z,q),q)}$
and therefore
(4.5) $\vartheta_{1}(y(z,q),z,q)_{\infty}=-1=y(z,q)M(Q(z,q),q).$
Inserting (4.4) into $Q^{\beta}(z,y)$ from (4.3) yields
$Q^{\beta}(z,y(z,q))=z^{\beta}M(Q(z,q),q)^{\beta\cdot E}=z^{\beta}\exp\big{(}{-a(-z(Q(z,q),q),q)}\big{)}^{\beta\cdot E}=Q^{\beta}(z,q).$
We see from Proposition 3.8 and the definition of $Q^{\beta}(z,y(z,q))$ in (4.3) that (4.5) takes the form $\theta_{1}(\widetilde{y})=\theta_{2}(\widetilde{y})$ for $\widetilde{y}=y(z,q)$. The final part of the statement of Proposition 4.1 now gives $\theta_{1}(y)=\theta_{2}(y)$ as elements of $A\llbracket y^{-1}\rrbracket[y]$ which gives the assertion. ∎
Corollary 4.3.
Put $q=e^{i\hbar}$. Then
$M(Q,q)=1+\sum_{\beta}\sum_{g\geq 0}\frac{1}{\beta\cdot E-1}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}.$
Proof.
The statement follows from Proposition 3.8 and Theorem 4.2. ∎
4.1. An equation satisfied by the full log potential
Definition 4.4.
We define the full log potential as
$W=\sum_{n>0}\vartheta_{n}U^{n}=\sum_{n>0}\sum_{m>0}\sum_{\beta:\beta\cdot{E}=n+m}\sum_{g\geq 0}mR_{g,(m,n)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}y^{-m}U^{n}.$
By [GRZ], Theorem 4.14, $W$ is a generating series of $2$-marked logarithmic Gromov-Witten invariants. However, it is not a generating series for open Gromov-Witten invariants, even in genus 00, because the open-log correspondence receives correction terms in both, higher genus and higher winding (resp. order of the fixed point). For genus 00 open Gromov-Witten invariants, the full potential can be explicitly written in terms of the winding $1$ invariants, [LM], (2.23). The aim of this section is to find an equation satisfied by the full log potential $W$. Since the coefficients of $\vartheta_{n}$ depend on $\vartheta_{1}$ in a difficult way, involving $2$-marked log invariants $R_{g,(p,q)}(\mathbb{P}(\log E),\beta)$ that are not known to have a simple generating function, this equation will be less elegant than the Lerche-Mayr equation. For a formal Laurent series $f(y)$, we use $\operatorname{Res}_{y=0}f(y)$ to refer to its formal residue, the $-1$st coefficient.
Proposition 4.5.
The full log potential $W$ satisfies the equation
$0=\operatorname{Res}_{y^{\prime}=0}W(y^{\prime},U)+\left(\frac{1}{U}-\vartheta_{1}(y)+\operatorname{Res}_{y^{\prime}=0}\left(D_{U}^{-1}W(y^{\prime},U)\right)\right)W(y,U).$
Proof.
By [GRZ], Proposition 5.2, the multiplication rule in the ring of theta functions gives
(4.6) $\vartheta_{n}\cdot\vartheta_{1}=\vartheta_{n+1}+\sum_{m=0}^{n}\alpha_{n,1}^{m}\vartheta_{m}+\alpha_{1,n}^{0},$
with
$\alpha_{n,1}^{m}=\sum_{\beta:\beta\cdot{E}=n-m+1}\sum_{g\geq 0}(n-m)R_{g,(n-m,1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta},$
and
$\alpha_{1,n}^{0}=\sum_{\beta:\beta\cdot{E}=n+1}\sum_{g\geq 0}R_{g,(1,n)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}.$
By inserting Proposition 3.8 into equation (4.6), we obtain a system of equations for the 2-marked invariants $R_{g,(m,n)}(\mathbb{P}(\log E),\beta)$. This system of equations determines all invariants $R_{g,(m,n)}(\mathbb{P}(\log E),\beta)$ from the invariants $R_{g,(n,1)}(\mathbb{P}(\log E),\beta)$, hence all theta functions $\vartheta_{n}(q)$ from the primitive one $\vartheta_{1}(q)$. To phrase the system of equations as an equation for the generating function $W$, we sum equation (4.6) in the sum $\sum_{n>0}(-)U^{n}$. The left hand side becomes $W\vartheta_{1}$ and the first term of the right hand side is $\frac{1}{U}W$. The last term on the right hand side is
$\operatorname{Res}_{y=0}W$ $=$ $\operatorname{Res}_{y=0}\sum_{n\geq 0}\sum_{m>0}\sum_{\beta:\beta\cdot{E}=n+m}\sum_{g\geq 0}mR_{g,(m,n)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}y^{-m}U^{n}$
$=$ $\sum_{n\geq 0}\sum_{\beta:\beta\cdot{E}=n+1}\sum_{g\geq 0}R_{g,(1,n)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}U^{n}.$
Similarly, the middle term on the right hand side is (where $D_{U}^{-1}$ is log integration, i.e., the inverse to $D_{U}=U\frac{d}{dU}$)
$\operatorname{Res}_{y^{\prime}=0}\left(D_{U}^{-1}W(y^{\prime},U)\right)W(y,U)=\operatorname{Res}_{y^{\prime}=0}\left(\sum_{m^{\prime}\geq 0}\frac{1}{m^{\prime}}\left((y^{\prime})^{m^{\prime}}\right.\right.$
$\qquad\qquad\left.\left.+\sum_{k>0}\sum_{\beta:\beta\cdot{E}=m^{\prime}+k}\sum_{g\geq 0}R_{g,(k,m^{\prime})}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}(y^{\prime})^{-k}\right)U^{m^{\prime}}\right)W(y,U)$
$=\left(\sum_{m^{\prime}\geq 0}\frac{1}{m^{\prime}}\sum_{\beta:\beta\cdot{E}=m^{\prime}+1}\sum_{g\geq 0}R_{g,(1,m^{\prime})}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}U^{m^{\prime}}\right)\left(\sum_{m\geq 0}\vartheta_{m}U^{m}\right)$
$=\sum_{n\geq 0}\sum_{m\geq 0}\sum_{\beta:\beta\cdot{E}=n-m+1}\sum_{g\geq 0}\frac{1}{n-m}R_{g,(1,n-m)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}\vartheta_{m}U^{n}$
$=\sum_{n\geq 0}\sum_{m\geq 0}\sum_{\beta:\beta\cdot{E}=n-m+1}\sum_{g\geq 0}(n-m)R_{g,(n-m,1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}z^{\beta}\vartheta_{m}U^{n}.$
Putting everything together, we obtain the claimed equation. ∎
5. Open-log correspondence for smooth log Calabi-Yau pairs
In this section, we prove a higher genus open-log correspondence for smooth log Calabi-Yau pairs $(\mathbb{P},E)$. We first show a correspondence between higher genus two-pointed log invariants and local invariants in a similar approach as in [GRZ]. Then assuming $\mathbb{P}$ is toric, we use the Topological Vertex [AKMV] to relate the log invariants to open invariants of $K_{\mathbb{P}}$ in all genera.
In order to state our results, we first establish some notation. Let $w\in\mathbb{Z}_{\geq 0}$, and $\vec{k}=(k_{1},\ldots,k_{n})\in\mathbb{Z}_{\geq 0}^{n}$ be a vector of non-negative integers such that $\sum_{j=1}^{n}jk_{j}=w$. Let $|\vec{k}|$ be the number of non-zero entries of $\vec{k}$. We denote $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be the genus-$g$, winding-$w$, open Gromov-Witten invariant of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}}$ in framing-0, in curve class $\beta+w\beta_{0}$ with $\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$, and in winding profile $\vec{k}$ if it has $k_{j}$ boundary components of winding $j$. The invariants $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ can be defined using stable relative maps, and we refer to [LLLZ] [FL] for a more detailed definition. We write $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to be the corresponding open-BPS invariant defined from open multiple cover formulas for $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ [MV]. We will frequently consider winding-$w$ invariants with 1 boundary component in which case the winding multiplicity vector is $\vec{k}=(0,\ldots,1,0,\ldots)$ has a single 1 in the $w$-th position corresponding to a single hole with winding $w$. As Gromov-Witten invariants and BPS invariants are uniquely determined by each other via the multiple-cover formulas, the $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ can be defined by the $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$.
Let $\pi:\hat{\mathbb{P}}\rightarrow\mathbb{P}$ be the blow up of $\mathbb{P}$ at a point with exceptional curve $C$. In Section 5.1, we first review the genus-0 open-log correspondence for smooth log Calabi-Yau pairs $(\mathbb{P},E)$ established in [GRZ]. In Section 5.2, we relate higher genus two-pointed log invariants $R_{g,(1,p)}(\mathbb{P}(\log E),\beta)$ with Gopakumar-Vafa invariants of local $\widehat{\mathbb{P}}$ in Theorem 5.2. In Section 5.3, additionally assuming $\mathbb{P}$ is toric, we use the Topological Vertex [AKMV] to relate $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ to closed invariants of $K_{\widehat{\mathbb{P}}}$ in curve class $\pi^{*}\beta-wC$ in Theorem 5.3, and then specialize it to winding-1 in Corollary 5.4 to obtain an open-closed equality of BPS invariants that generalizes Theorem 1.1, [LLW] to all genera. In Section 5.5, we prove in Theorem 5.8 an open-log correspondence in all genera that relates two-pointed log invariants $R_{g,(1,p)}(\mathbb{P}(\log E),\beta)$ to winding-1, 1-boundary component, open-BPS invariants $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ via a discrepancy quantified by stationary Gromov-Witten invariants of $E$, and $R_{h,(1,p)}(\mathbb{P}(\log E),\beta)$ for $h<g$. We provide a computational check of Theorem 5.8 in Example 5.10 for $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$. In Section 5.6, we provide a computational check of Corollary 5.4 using the topological vertex with the relevant open and closed BPS invariants listed in Table 1.
5.1. The $g=0$ open-log correspondence
We have the following genus-0 open-log correspondence due to [GRZ].
Theorem 5.1 (Theorem 1.1, [GRZ]).
Let $\mathbb{P}$ be a toric Fano surface with smooth anticanonical divisor $E$. Let $\beta\in\textup{NE}(\mathbb{P})$ be an effective curve class and $\beta_{0}\in H_{2}(K_{\mathbb{P}},L)$ be a holomorphic disc class with boundary on $L$. Then,
$O_{0}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))=(-1)^{\beta\cdot E}(\beta\cdot E-1)R_{0,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$
When $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$, the above invariants are the coefficients of the open mirror map $M(Q)$ of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{C}\mathbb{P}^{2}}$ in framing-0,
$M(Q)=1-2Q+5Q^{2}-32Q^{3}+286Q^{4}-3038Q^{5}+\ldots$
Theorem 5.1 implies that the mirror dual Landau-Ginzburg superpotential is equivalent to the open mirror map $M(Q)$. Its proof uses results from [Gra2], [vGGR], [LLW], and Corollary 6.6 of [GRZ] to relate two-pointed log invariants $R_{0,(1,\beta\cdot E-1)}$ to open invariants of $K_{\mathbb{P}}$.
5.2. Higher genus log/Gopakumar-Vafa correspondence
Let $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ be the genus-$g$, local Gromov-Witten invariant of $K_{\widehat{\mathbb{P}}}$ in class $\pi^{*}\beta-C$ [CKYZ], and $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ the genus-$g$, Gopakumar-Vafa invariant of $K_{\widehat{\mathbb{P}}}$ in curve class $\pi^{*}\beta-C$ defined by multiple cover formulas [GV1] [GV2]. Let $Q^{\beta}$ be a formal variable tracking the curve class $\beta\in\textup{NE}(\mathbb{P})$. Using the higher genus analogues of arguments used to prove Theorem 5.1, we have the following theorem.
Theorem 5.2.
Let $(\mathbb{P},E)$ be a smooth log Calabi-Yau pair. Then,
$\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\frac{1}{(\beta\cdot E-1)}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}=\\ \sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\left[(-1)^{\beta\cdot E}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\left(2\sin\frac{\hbar}{2}\right)^{2g-2}Q^{\beta}\right]-\Delta$
where $\Delta$, defined in Equation 5.2, is a discrepancy expressed by the stationary Gromov-Witten theory of $E$ and two pointed log invariants of $\mathbb{P}(\log E)$.
Proof of Theorem 5.2.
From here on, we simplify the notation by indexing sums ${\sum_{g,\beta}}$ where there is no confusion. We take the generating series of two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(X(\log E),\beta)$ with one prescribed contact point with $E$ of order 1 and one non-prescribed contact point with $E$ of order $(\beta\cdot E-1)$,
$\sum_{g,\beta}\frac{1}{(\beta\cdot E-1)}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}$
By Corollary 6.6 of [GRZ], the above generating series becomes,
(5.1) $\begin{split}\sum_{g,\beta}\frac{1}{(\beta\cdot E-1)}&\left[R_{g}(\widehat{\mathbb{P}}(\log\pi^{*}E-C),\pi^{*}\beta-C)\right.\\ &-\left.\sum_{i=0}^{g-1}R_{i,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)N(g-i,1)\right]\hbar^{2g}Q^{\beta}\\ \end{split}$
where $R_{g}(\widehat{\mathbb{P}}(\log\pi^{*}E-C),\pi^{*}\beta-C)$ are genus-$g$, maximal tangency invariants of the log Calabi-Yau pair (^$(\widehat{\mathbb{P}},\pi^{*}E-C)$ with $\lambda_{g}$-insertion (defined in Section 4, [GRZ]), and $N(g,p)$ is the genus-$g$, local relative invariant of $\mathbb{P}^{1}(\log\infty)$ with tangency order $p$ computed in Theorem 5.1, [BP], and is given by the $\hbar^{2g}$-coefficient of $\frac{(-1)^{p+1}}{p}\frac{i\hbar}{q^{p/2}-q^{-p/2}}$.
For reasons we shall see, define $\Delta$ to be the term,
(5.2) $\begin{split}\Delta:=\sum_{g,\beta}\Biggl{[}&(-1)^{\beta\cdot E}\biggl{[}\sum_{n\geq 0}\sum_{\begin{subarray}{c}g=h+g_{1}+\ldots+g_{n},\\ \mathbf{a}=(a_{1},\ldots,a_{n})\in\mathbb{Z}^{n}_{\geq 0},\\ \pi^{*}\beta-C=d_{E}[E]+\beta_{1}+\ldots+\beta_{n},\\ d_{E}\geq 0,\beta_{j}\cdot D>0\end{subarray}}\frac{(-1)^{g-1+(E\cdot E)d_{E}}(E\cdot E)^{m}}{m!|Aut(\mathbf{a},g)|}\\ &N_{h,(\mathbf{a},1^{m})}(E,d_{E})\prod_{j=1}^{n}((-1)^{\beta_{j}\cdot E}(\beta_{j}\cdot E)R_{g_{j},(\beta_{j}\cdot E)}(\widehat{\mathbb{P}}(\log\pi^{*}E-C),\beta_{j})\biggr{]}\\ &-\frac{1}{(\beta\cdot E-1)}\sum_{i=0}^{g-1}R_{i,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)N(g-i,1)\Biggr{]}\hbar^{2g}Q^{\beta}\end{split}$
[…]
(5.3) $N_{h,(\textbf{a},1^{m})}(E,d_{E}):=\int_{[\overline{\mathcal{M}}_{h,n+m}(E,d_{E})]^{vir}}\prod_{i=1}^{n}ev_{i}^{*}[pt]\psi_{i}^{a_{i}}\prod_{j=1}^{m}ev_{j}^{*}[pt]\psi_{j}$
where $\textbf{a}\in\mathbb{Z}^{n}_{\geq 0}$, $\psi_{j}\in H^{2}(\overline{\mathcal{M}}_{h,n+m}(E,d_{E}))$ is the $\psi-$class at the $j$-th marked point, and $[pt]\in H^{2}(E)$ is the Poincare-dual of a point.
For $g\geq 0$ and $\beta\in\textup{NE}(\mathbb{P})$, we define $\Delta(g,\beta)$ as the $\hbar^{2g}Q^{\beta}$-coefficient of $\triangle$ by,
$\Delta:=\sum_{g,\beta}\Delta(g,\beta)\hbar^{2g}Q^{\beta}$
Applying the $g\geq 0$ log-local principle (combining Propositions 3.1 and 3.4 of [BFGW]) to $R_{g,(\beta\cdot E-1)}(\widehat{\mathbb{P}}(\log\pi^{*}E-C),\pi^{*}\beta-C)$ in (5.1), it becomes,
(5.4) $\begin{split}\sum_{g,\beta}\left[(-1)^{\beta\cdot E}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g}Q^{\beta}\right]-\Delta\end{split}$
As $\pi^{*}\beta-C$ is a primitive curve class, the Gopakumar-Vafa formula for $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ takes the form,
(5.5) $\sum_{g\geq 0}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}=\sum_{g\geq 0}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\left(2\sin\frac{\hbar}{2}\right)^{2g-2}$
By combining (5.1), (5.2), (5.4), and (5.5), we have,
(5.6) $\begin{split}&\sum_{g,\beta}\frac{1}{(\beta\cdot E-1)}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}=\\ &\hskip 21.68121pt\sum_{g,\beta}\left[(-1)^{\beta\cdot E}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\left(2\sin\frac{\hbar}{2}\right)^{2g-2}Q^{\beta}\right]-\Delta\\ \end{split}$
Theorem 5.2 relates the generating function of genus-$g$, two-pointed log invariants $R_{g,(\beta\cdot E-1,1)}(\mathbb{P}(\log E),\beta)$ to the generating function of genus-$g$, Gopakumar-Vafa invariants $n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ in curve class $\pi^{*}\beta-C$. We give an explicit expression for the discrepancy $\Delta$ in Example 5.10.
5.3. Winding-$w$ open/closed correspondence
For this section, suppose that $\mathbb{P}$ is toric and $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up at a point. Then $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ are toric Calabi-Yau 3-folds. We use the Topological Vertex [AKMV] to derive an equality relating the winding-$w$, open invariants $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$ of an outer AV-brane $L\subset K_{\mathbb{P}}$ in winding profile $\vec{k}$ to the closed invariants $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$. The work of [LLLZ] [LLZ1] [LLZ2] [MOOP] show that calculations with the Topological Vertex compute the open Gromov-Witten invariants $O_{g}(K_{\mathbb{P}}/L,\beta+w\beta_{0},\vec{k})$. We refer to [Kon] [LLLZ] [AKMV] for more definitions and details on the topological vertex.
The toric diagram of a local toric Fano surface $\mathbb{P}$ is given by a convex $k$-gon with a non-compact edge emanating from each vertex. We refer to Figure 5.1 for the relevant toric diagrams. Let $\{v_{i}\}_{i=1}^{k}$ be the torus-fixed points, and $\{\ell_{i}\}_{i=1}^{k}$ be the compact 1-dimensional torus orbits of the toric diagram of local $\mathbb{P}$. Let $t_{i}$ be the Kähler parameter of $\ell_{i}$, and define the variable $Q^{\ell_{i}}:=e^{-t_{i}}$. For each $\ell_{i}$, we assign a variable $\lambda_{i}$ that runs over all partitions of arbitrary length. We refer to a partition $\lambda$ and its $U(N$)-representation interchangeably. By the topological vertex, the partition function of local $\mathbb{P}$ is given by,
$Z_{\mathbb{P}}(Q^{\ell_{i}})=\sum_{\lambda_{1},\ldots,\lambda_{k}}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\left(\prod_{j=1}^{k-1}C_{\lambda^{T}_{j}\lambda_{j+1}\varnothing}\right)C_{\lambda_{k}^{T}\lambda_{1}\varnothing}$
where the vertex functions $C_{\lambda_{1}\lambda_{2}\lambda_{3}}$ are defined in Definition 2.2, [Kon].
The toric diagram of local $\widehat{\mathbb{P}}$ has an additional compact 1-dimensional torus orbit in the polygon corresponding to the exceptional curve $C\cong\mathbb{P}^{1}$, with Kähler parameter $t_{k+1}$ and $Q^{C}:=e^{-t_{k+1}}$. We label the partition of $C$ to be $\lambda_{k+1}$. The partition function of local $\widehat{\mathbb{P}}$ is given by,
$Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{C})=\sum_{\lambda_{1},\ldots,\lambda_{k+1}}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}Q^{|\lambda_{k+1}|C}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\left(\prod_{j=1}^{k}C_{\lambda^{T}_{j}\lambda_{j+1}\varnothing}\right)C_{\lambda_{k+1}^{T}\lambda_{1}\varnothing}$
In addition, we define $Z_{\widehat{\mathbb{P}}}|_{\pi^{*}\beta-wC}$ to be,
$Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{C})|_{\pi^{*}\beta-wC}:=Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{-C})|_{|\lambda_{k+1}|=w}$
As the normal bundle of $C$ in $K_{\widehat{\mathbb{P}}}$ is $\mathcal{O}_{\mathbb{P}^{1}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{1}}(-1)$, denote $W$ to be the space obtained by flopping along $C\subset K_{\widehat{\mathbb{P}}}$. We write $E$ for the exceptional curve of $W$, with Kähler parameter $t^{\prime}_{k+1}$ and $Q^{E}:=e^{-t^{\prime}_{k+1}}$. We label the partition of $E$ to be $\lambda^{\prime}_{k+1}$. The partition function of $W$ is given by,
$Z_{W}(Q^{\ell_{i}},Q^{E})=\sum_{\lambda_{1},\ldots,\lambda_{k},\lambda^{\prime}_{k+1}}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}Q^{|\lambda^{\prime}_{k+1}|E}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\left(\prod_{j=1}^{k-1}C_{\lambda^{T}_{j}\lambda_{j+1}\varnothing}\right)C_{\lambda^{T}_{k}\lambda_{1}\lambda^{\prime}_{k+1}}C_{\lambda^{{}^{\prime}T}_{k+1}\varnothing\varnothing}$
$C$$E$$v_{k}$$v_{1}$$v_{k}$$v_{k+1}$$v_{1}$$\lambda_{k+1}$$v_{k+1}$$\lambda^{\prime}_{k+1}$$\lambda_{1}$$\lambda_{1}$$\lambda_{k}$$v_{k-1}$$\lambda_{k}$$v_{k-1}$$\widehat{\mathbb{P}}$$\mathbb{P}$$v_{k}$$v_{1}$$\lambda_{1}$$v_{k-1}$$\lambda_{k}$
Figure 5.1. Toric diagrams of ^$\widehat{\mathbb{P}},W$, and $\mathbb{P}$.
Let $V$ be an $U(N)$-matrix with $N$ suitably large, representing the monodromy of a flat connection on $L$. Let $Tr_{\nu}V$ be the trace of $V$ in a $U(N)$-representation given by a partition $\nu$. The open free energy of a single outer AV-brane $L\subset K_{\mathbb{P}}$ is given by,
$Z_{\mathbb{P}}(Q^{\ell_{i}},V)=\sum_{\lambda_{1},\ldots,\lambda_{k},\nu}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\left(\prod_{j=1}^{k-1}C_{\lambda^{T}_{j}\lambda_{j+1}\varnothing}\right)C_{\lambda^{T}_{k}\lambda_{1}\nu}Tr_{\nu}V$
For a fixed partition or representation $\nu$ attached to an outer AV-brane, we define,
$Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\nu}:=\sum_{\lambda_{1},\ldots,\lambda_{k}}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\left(\prod_{j=1}^{k-1}C_{\lambda^{T}_{j}\lambda_{j+1}\varnothing}\right)C_{\lambda^{T}_{k}\lambda_{1}\nu}Tr_{\nu}V$
We have the following theorem.
Theorem 5.3.
Let $\mathbb{P}$ be a toric del Pezzo surface, and $\beta\in\textup{NE}(\mathbb{P})$ an effective curve class. Let $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ be the toric blow up at a point with exceptional curve $C$. Let $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ be the toric canonical bundles. We have the equality of winding-$w$ open and closed partition functions,
$\sum_{|\lambda|=w}Z_{\mathbb{P}}(V)|_{\lambda}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{T}}=Z_{\widehat{\mathbb{P}}}|_{\pi^{*}\beta-wC}$
where $Z_{\mathbb{C}^{3}}(V)|_{\lambda^{T}}$ is the winding-$w$, open partition function of an outer AV-brane in $\mathbb{C}^{3}$ in partition $\lambda^{T}$.
Equivalently, writing $\beta=c_{1}\ell_{1}+\ldots c_{k}\ell_{k}$, the invariant $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ is the $\hbar^{2g}$-coefficient of the expression,
$\sum_{\begin{subarray}{c}l\geq 0,l|w,\\ (w_{1},\ldots,w_{l})\in\mathbb{Z}^{l}_{\geq 0},\\ 1\leq j\leq l,|(\lambda^{\prime}_{k+1})_{j}|=w_{j}\end{subarray}}\prod_{i=1}^{l}\sum_{\begin{subarray}{c}1\leq a\leq k,\\ |\lambda_{ia}|=c_{ia},\\ c_{ia}\in\mathbb{Z}_{\geq 0}\end{subarray}}\left[\frac{(-1)^{l+1+\sum_{a}|\lambda_{ia}|}}{l}q^{\sum_{a}\kappa_{\lambda_{ia}}}\left(\prod_{a=1}^{k-1}C_{\lambda^{T}_{ia}\lambda_{(i+1)a}\varnothing}\right)C_{\lambda^{T}_{ka}\lambda_{1a}(\lambda^{\prime}_{k+1})_{i}}C_{(\lambda^{{}^{\prime}T}_{k+1})_{i}\varnothing\varnothing}\right]$
where $q=e^{i\hbar}$, $(w_{1},\ldots,w_{l})\in\mathbb{Z}_{\geq 0}^{l}$ is a partition of $w$, $(\lambda^{\prime}_{k+1})_{j}$ is a partition of size $w_{j}$ for $1\leq j\leq l$, $(c_{ia})\in\mathbb{Z}_{\geq 0}^{l}$ is a partition of $c_{a}$ for $1\leq a\leq k$.
Proof of Theorem 5.3.
By the gluing formula for the topological vertex or Proposition 5.8, [LLLZ], which is shown by the degeneration formula for relative Gromov-Witten invariants [Li], we can express $Z_{W}$ in terms of winding-$w$, open Gromov-Witten invariants of an outer AV-brane $L\subset K_{\mathbb{P}}$ and the open invariants of an outer AV-brane in $\mathbb{C}^{3}$. We have,
$Z_{W}(Q^{\ell_{i}},Q^{E})$ $=\sum_{\lambda^{\prime}_{k+1}}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}$
where $Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}$ is the partition function of winding-$w$, open Gromov-Witten invariants of an outer AV-brane in $K_{\mathbb{P}}$ in representation $\lambda^{\prime}_{k+1}$, and $Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}$ is the partition function of winding-$w$, open Gromov-Witten invariants of an outer AV-brane in $\mathbb{C}^{3}$ in representation $\lambda^{{}^{\prime}T}_{k+1}$. In particular, from Equation (6.9), [BBvG], we have,
$Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}:=C_{\lambda^{\prime}_{k+1}\varnothing\varnothing}=q^{\kappa_{\lambda^{\prime}_{k+1}}/2}\sum_{\eta}s_{\lambda^{{}^{\prime}T}_{k+1}/\eta}(q^{\rho})$
By flop invariance of the topological vertex [KM], we have,
(5.7) $Z_{W}(Q^{\ell_{i}},Q^{E})=Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{-C})$
via the equality of Kähler parameters $Q^{E}=Q^{-C}$. Equivalently, Equation 5.7 is,
(5.8) $\begin{split}&\sum_{\lambda^{\prime}_{k+1}}Q^{|\lambda^{\prime}_{k+1}|E}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}\\ &=\sum_{\lambda_{k+1}}Q^{-|\lambda_{k+1}|C}\sum_{\lambda_{1},\ldots,\lambda_{k}}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\left(\prod_{j=1}^{k}C_{\lambda^{T}_{j}\lambda_{j+1}\varnothing}\right)C_{\lambda_{k+1}^{T}\lambda_{1}\varnothing}\end{split}$
Restricting to $|\lambda^{\prime}_{k+1}|=w$, the coefficient of $Q^{wE}$ must equal the coefficient of $Q^{-wC}$. Therefore, we have the first statement in Theorem 5.3,
$\sum_{|\lambda^{\prime}_{k+1}|=w}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}=Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{C})|_{\pi^{*}\beta-wC}$
where $Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{C})|_{\pi^{*}\beta-wC}$ is defined in Section 5.3.
For the second statement in Theorem 5.3, we take the logarithm of the partition function to get the free energy, i.e., the generating function of connected Gromov-Witten invariants. We have,
$\log Z_{W}(Q^{\ell_{i}},Q^{E})$ $=\log\left(1+\sum_{\lambda^{\prime}_{k+1},\exists\lambda_{i}\neq 0}Q^{|\lambda^{\prime}_{k+1}|E}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}\right)$
$=\sum_{g,\beta,w\in\mathbb{Z}_{\geq 0}}N_{g}(W,\beta+wE)\hbar^{2g}Q^{\beta+wE}$
and,
$\log Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{-C})$
$=\log\left(1+\sum_{\exists\lambda_{i}\neq 0}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}Q^{-|\lambda_{k+1}|C}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\left(\prod_{j=1}^{k}C_{\lambda^{T}_{j}\lambda_{(j+1)}\varnothing}\right)\right.\left.C_{\lambda_{k+1}^{T}\lambda_{1}\varnothing}\right)$
$=\sum_{g,\beta,w\in\mathbb{Z}_{\geq 0}}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)\hbar^{2g}Q^{\beta-wC}$
The equality of partition functions $\log Z_{W}=\log Z_{\widehat{\mathbb{P}}}$ implies that the coefficient of $Q^{-|\lambda_{k+1}|C}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}$ in $\log Z_{\widehat{\mathbb{P}}}$ is equal to the coefficient of $Q^{|\lambda^{\prime}_{k+1}|E}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}$ in $\log Z_{W}$, i.e., $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)=N_{g}(W,\beta+wE)$.
We express $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ explicitly in terms of vertex functions of $Z_{W}$. We define $S:=\sum_{\begin{subarray}{c}\lambda^{\prime}_{k+1},\\ \exists\lambda_{i}\neq 0\end{subarray}}Q^{|\lambda^{\prime}_{k+1}|E}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}$, hence $\log Z_{W}=\log(1+S)=\sum_{l\geq 0}\frac{(-1)^{l+1}}{l}S^{l}$.
We find the coefficient of $Q^{\beta+wE}$ in $\log Z_{W}$. We only consider those powers $S^{l}$ where $l|w$. Let $(w_{1},\ldots,w_{l})$ be a partition of $w$ where $\sum_{i}w_{i}=w$ with $w_{i}\geq 0$. We see that the coefficient of $Q^{wE}$ in $S^{l}$ is,
$\sum_{(w_{1},\ldots,w_{l})}\sum_{\begin{subarray}{c}1\leq j\leq l,\\ |(\lambda^{\prime}_{k+1})_{j}|=w_{j}\end{subarray}}P_{1}\ldots P_{l}$
where we define $P_{1}\ldots P_{l}:=\left(\prod_{i=1}^{l}Z_{\mathbb{P}}(V)|_{(\lambda^{\prime}_{k+1})_{i}}Z_{\mathbb{C}^{3}}(V)|_{(\lambda^{{}^{\prime}T}_{k+1})_{i}}\right)$, and $(\lambda^{\prime}_{k+1})_{j}$ is the representation attached to the external leg for the $j$-th term in $P_{j}$ for $1\leq j\leq l.$
Suppose we write $\beta=c_{1}\ell_{1}+\ldots c_{k}\ell_{k}$ in the basis given by the compact 1-dimensional torus orbits. Let $(\beta_{1},\ldots,\beta_{l})$ be a partition of $\beta$, i.e., $\sum_{i=1}^{l}\beta_{i}=\beta$, and we write $\beta_{i}=c_{i1}\ell_{1}+\ldots+c_{ik}\ell_{k}$ with $\sum_{i=1}^{l}c_{ia}=c_{a}$ for $a=1,\ldots,k$, and $c_{ia}\geq 0$. For $1\leq i\leq l$, we suppose that the coefficient of $Q^{\beta_{i}}$ appears only in $P_{i}$. It is given by,
$\sum_{\begin{subarray}{c}1\leq a\leq k,\\ |\lambda_{ia}|=c_{ia}\end{subarray}}(-1)^{\sum_{a}|\lambda_{ia}|}q^{\sum_{a}\kappa_{\lambda_{ia}}}\left(\prod_{a=1}^{k-1}C_{\lambda^{T}_{ia}\lambda_{(i+1)a}\varnothing}\right)C_{\lambda^{T}_{ka}\lambda_{1a}(\lambda^{\prime}_{k+1})_{i}}C_{(\lambda^{{}^{\prime}T}_{k+1})_{i}\varnothing\varnothing}$
where $\lambda_{ia}$ for $a=1,\ldots,k$ are the partitions used in $P_{i}$. Taking the product over all $i$ gives the coefficient of $Q^{\beta}$ in $S^{l}$. Thus, also summing over $l$, the coefficient of $Q^{\beta+wE}$ is,
$\sum_{\begin{subarray}{c}l\geq 0,l|w,\\ (w_{1},\ldots,w_{l})\in\mathbb{Z}^{l}_{\geq 0},\\ 1\leq j\leq l,|(\lambda^{\prime}_{k+1})_{j}|=w_{j}\end{subarray}}\prod_{i=1}^{l}\sum_{\begin{subarray}{c}1\leq a\leq k,\\ |\lambda_{ia}|=c_{ia},\\ c_{ia}\in\mathbb{Z}_{\geq 0}\end{subarray}}\left[\frac{(-1)^{l+1+\sum_{a}|\lambda_{ia}|}}{l}q^{\sum_{a}\kappa_{\lambda_{ia}}}\left(\prod_{a=1}^{k-1}C_{\lambda^{T}_{ia}\lambda_{(i+1)a}\varnothing}\right)C_{\lambda^{T}_{ka}\lambda_{1a}(\lambda^{\prime}_{k+1})_{i}}C_{(\lambda^{{}^{\prime}T}_{k+1})_{i}\varnothing\varnothing}\right]$
Substituting $q=e^{i\hbar}$ in the above, the resulting $\hbar^{2g}$-coefficient is the invariant $N_{g}(W,\beta+wE)$, which is equal to $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ by flop invariance [KM]. ∎
Theorem 5.3 expresses the closed invariant $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ of local $\widehat{\mathbb{P}}$ in terms of a product of open invariants of an outer AV-brane in local $\mathbb{P}$ and open invariants of an outer AV-brane in $\mathbb{C}^{3}.$
5.4. Application: winding-1 open-closed equality
Specializing Theorem 5.3 to winding-1 open invariants, we have the following corollary equating open and closed BPS invariants in all genera.
Corollary 5.4.
Suppose $\mathbb{P}$ is toric, and $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up at a point with exceptional curve $C$. Let $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ be the toric canonical bundles. Then, we have
$n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)=(-1)^{g+1}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$
For some explicit values of open and closed BPS invariants, we refer to Section 5, [GZ] and Appendix B, [HKR] respectively.
Proof of Corollary 5.4.
We first derive a statement equivalent to Corollary 5.4. The Gopakumar-Vafa formula [GV1] [GV2] for $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ states,
(5.9) $\sum_{g\geq 0}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}=\sum_{g\geq 0}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\left(2\sin\frac{\hbar}{2}\right)^{2g-2}$
The open multiple cover formula [MV] for $O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ states,
(5.10) $\sum_{g\geq 0}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}=\sum_{g\geq 0}(-1)^{g+1}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\left(2\sin\frac{\hbar}{2}\right)^{2g-1}$
From Equation 5.10, Corollary 5.4 is equivalent to the equality,
(5.11) $\sum_{g\geq 0}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}=\sum_{g\geq 0}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\left(2\sin\frac{\hbar}{2}\right)^{2g-1}$
After dividing by $\left(2\sin\frac{\hbar}{2}\right)$, plugging in Equation 5.9, and summing over $\beta\in H_{2}^{+}(\mathbb{P},\mathbb{Z})$ in Equation 5.11, we have the following equality,
(5.12) $\sum_{g\geq 0}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}Q^{\beta}=\left(2\sin\frac{\hbar}{2}\right)\sum_{g\geq 0}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}Q^{\beta}$
Therefore, it suffices to show that Equation 5.12 holds. Specializing to winding-1 open invariants, the first statement of Theorem 5.3 gives the equality,
$Z_{\widehat{\mathbb{P}}}|_{\pi^{*}\beta-C}=Z_{\mathbb{P}}(V)|_{\square}Z_{\mathbb{C}^{3}}(V)|_{\square}$
This is equivalently expressed in Equation 5.8 as,
$Q^{-C}\sum_{\lambda_{1},\ldots,\lambda_{k}}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\prod_{j=1}^{k+1}C_{\lambda^{T}_{j}\lambda_{(j+1)\text{mod }k+1}\varnothing}=Q^{E}Z_{\mathbb{P}}(V)|_{\square}Z_{\mathbb{C}^{3}}(V)|_{\square}$
Since the curve class $\pi^{*}\beta-C$ is primitive, we obtain the winding-1 invariants directly from the partition functions. Hence we have,
$Q^{-C}\sum_{g,\beta}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}Q^{\beta}=Q^{E}\left(\sum_{g,\beta}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}Q^{\beta}\right)Z_{\mathbb{C}^{3}}(V)|_{\square}$
After canceling $Q^{E}$ with $Q^{-C}$ by flop invariance [KM] and noticing that $Z_{\mathbb{C}^{3}}(V)|_{\square}:=C_{\square\varnothing\varnothing}=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{-1}$ (from pg.53, [AKMV]), we have
$\sum_{g,\beta}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}Q^{\beta}=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})\sum_{g,\beta}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}Q^{\beta}$
The result follows after making the variable change $q=e^{i\hbar}$, and redefining the generating function $\sum_{g,\beta}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}Q^{\beta}$ up to a global factor of $i$. ∎
Remark 5.5.
In genus-0, we have $O_{0}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))=-n_{0}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ and $N_{0}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)=n_{0}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ since $\pi^{*}\beta-C$ is a primitive curve class. By remarks made in Section 2.2 of [GRZ], $O_{0}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ equals the genus-0, open Gromov-Witten invariant of a moment fiber of $K_{\mathbb{P}}$. Thus, Corollary 5.4 is equivalent to Theorem 1.1, [LLW] in genus-0.
Remark 5.6.
For toric Fano manifolds, the genus-0, open Gromov-Witten invariant of a moment fiber was defined in the symplectic category by [FOOO]. As noted in Remark 5.5, both the open invariant of an outer AV-brane $L\subset K_{\mathbb{P}}$ or of a moment fiber agree in genus-0. By Corollary 5.4 and Theorem 1.1, [LLW], genus-0 open invariants defined either using stable relative maps as in [LLLZ] [FL] or in the symplectic category agree since they are both equal to $N_{0}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$.
Remark 5.7.
Corollary 5.4 is also used to establish a correspondence between the winding-1, open-BPS invariants $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ of an outer AV-brane $L\subset K_{\mathbb{P}}$ and certain higher genus closed invariants of the projectivized canonical bundle $\mathbb{P}(K_{\mathbb{P}}\oplus\mathcal{O}_{\mathbb{P}})$ [Zho].
5.5. Higher genus open-log correspondence
From previously shown results in Section 5, we have a higher genus open-log correspondence for smooth log Calabi-Yau pairs $(\mathbb{P},E)$ that relates the generating function of two-pointed log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ to the genus-$g$, winding-1, 1-boundary component, open-BPS invariants $n^{open}_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ of an outer AV-brane $L\subset K_{\mathbb{P}}$ in framing-0,
Theorem 5.8 (Open-log correspondence for smooth log Calabi-Yau pairs).
Assume that $\mathbb{P}$ is toric, and $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up. Then, we have,
$\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\frac{1}{(\beta\cdot E-1)}R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)\hbar^{2g}Q^{\beta}=\\ \sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in\textup{NE}(\mathbb{P})\end{subarray}}\left[(-1)^{\beta\cdot E+g+1}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\left(2\sin\frac{\hbar}{2}\right)^{2g-2}Q^{\beta}\right]-\Delta$
where the discrepancy term $\Delta$ is as in Theorem 5.2.
Proof.
Apply Corollary 5.4 to Theorem 5.2. ∎
Thus, by Theorem 5.8, we can express log invariants $R_{g,(1,\beta\cdot E-1)}(\mathbb{P}(\log E),\beta)$ by open-BPS invariants $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$, and vice versa.
Remark 5.9.
Open-log correspondences in all genera have been proven when the anti-canonical divisor is singular. In [BBvG], an open-log correspondence is proven for Looijenga pairs satisfying a deformation property, and is extended to quasi-tame Looijenga pairs in [BrSc]. In [Sch], an all genus open-log correspondence is established for two-component Looijenga pairs.
Example 5.10.
Let $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$. We may write $O_{1}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))=n_{1}^{open}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))-\frac{1}{24}n_{0}^{open}(K_{\mathbb{P}^{2}}/L,\beta+\beta_{0},(1))$ by the open multiple cover formulas [MV]. Specializing Theorem 5.8 to genus-1 and degree-4, we have the following equality,
$O_{1}(K_{\mathbb{C}\mathbb{P}^{2}}/L,4H+\beta_{0},(1))=\frac{1}{11}R_{1,(1,11)}(\mathbb{C}\mathbb{P}^{2}(\log E),4H)+\frac{11}{24}R_{0,(1,11)}(\mathbb{C}\mathbb{P}^{2}(\log E),4H)\\ -3R_{0,(3)}(\mathbb{F}_{1}(\log\pi^{*}E-C),\pi^{*}H)-2R_{0,(1)}(\mathbb{F}_{1}(\log\pi^{*}E-C),C)R_{0,(2)}(\mathbb{F}_{1}(\log\pi^{*}E-C),F)$
We can compute that,
$\Delta(1,4H)=O_{1}(K_{\mathbb{C}\mathbb{P}^{2}}/L,4H+\beta_{0},(1))-\frac{1}{11}R_{1,(1,11)}(\mathbb{C}\mathbb{P}^{2}(\log E),4H)+\frac{1}{24}n_{0}(K_{\mathbb{F}_{1}},\pi^{*}4H-C)$
where $\Delta(1,4H)$ is defined in Equation 5.2.
The open Gromov-Witten invariant $O_{1}(K_{\mathbb{C}\mathbb{P}^{2}}/L,4H+\beta_{0},(1))$ was computed by localization to be $\frac{-3313}{12}$ [GZ]. The relevant right hand side invariants $R_{g,(11,1)}(\mathbb{C}\mathbb{P}^{2}(\log E),4H)$ and $R_{0}(\mathbb{F}_{1}(\log\pi^{*}E-C),\beta)$ can be computed by $q$-refined tropical curve counting in the scattering diagram of $(\mathbb{C}\mathbb{P}^{2},E)$ ([GRZ], Appendix B) and/or using the formula of [vGGR]. The first equality above is explicitly,
$\frac{-3313}{12}=\frac{1}{11}\cdot(-18513)+\frac{11}{24}\cdot 3146-3\cdot 9-2\cdot 1\cdot 4.$
Remark 5.11.
Recall from Theorem 4.2 that $\vartheta_{1}(q)$ or equivalently $M(Q,q)$ encode the higher-genus, two-pointed log invariants of $(\mathbb{P},E)$. We note that it is not the generating series of higher-genus, winding-1, 1 boundary component, open invariants of an outer Aganagic-Vafa brane $L\subset K_{\mathbb{P}}$. As an example, consider $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$. We have from Example 2.27 that,
$M(Q,q)=1-(q^{\frac{1}{2}}+q^{\frac{-1}{2}})Q+(q^{-2}+q^{-1}+1+q+q^{2})Q^{2}\\ -\left(q^{\frac{9}{2}}+3q^{\frac{7}{2}}+4q^{\frac{5}{2}}+4q^{\frac{3}{2}}+4q^{\frac{1}{2}}+4q^{-\frac{1}{2}}+4q^{-\frac{3}{2}}+4q^{-\frac{5}{2}}+3q^{-\frac{7}{2}}+q^{-\frac{9}{2}}\right)Q^{3}+\ldots$
whereas the higher-genus open invariants of $L\subset K_{\mathbb{C}\mathbb{P}^{2}}$ can be computed by the Topological Vertex (Equation (9.10), [AKMV]),
$\widehat{f}_{\square\,\varnothing\,\varnothing}(z,Q)=1-2Q+5Q^{2}-(32+9z)Q^{3}+(286+288z+108z^{2}+14z^{3})Q^{4}+\ldots$
where $z=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2}$. The coefficient of $z^{g}Q^{d}$ in $\widehat{f}_{\square\,\varnothing\,\varnothing}$ is the negative of the genus-$g$, degree-$d$, winding-1, 1 boundary component, LMOV invariant of an outer AV-brane in local $\mathbb{C}\mathbb{P}^{2}.$ For an outer AV-brane, the winding-1, LMOV invariant is equal to the winding-1, open-BPS invariant $n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$.  [^6] In general, LMOV invariants are related to open-BPS invariants by a linear transformation defined by characters of the symmetric group [AKMV]. Note that in the limit as $z\rightarrow 0$ or $q\rightarrow 1$, we recover the open mirror map $M(Q,q)$ from $\widehat{f}_{\square\,\varnothing\,\varnothing}$. From Example 2.27, we see that $\widehat{f}_{\square\,\varnothing\,\varnothing}\neq M(Q,q)$, with the difference $M(Q,q)-\widehat{f}_{\square\,\varnothing\,\varnothing}$ effectively described by Theorem 5.8.
5.6. Verification of Corollary 5.4
We provide a computational check of Corollary 5.4 and Theorem 5.8 for $\mathbb{P}=\mathbb{C}\mathbb{P}^{2}$ for curve classes $\beta=dH$ for $d\leq 4$ and all genera. To perform calculations with the topological vertex, we use the diagrams in Figure 5.2. The winding-1, open topological string partition function $Z_{K_{\mathbb{C}\mathbb{P}^{2}}}(V)$ of an outer $AV$-brane in local $\mathbb{P}^{2}$ in framing-0 is given by,
$Z_{\mathbb{C}\mathbb{P}^{2}}(V)=\sum_{\lambda_{1},\lambda_{2},\lambda_{3}}(-1)^{\sum_{i}|\lambda_{i}|}Q^{\left(\sum_{i}|\lambda_{i}|\right)H}q^{\sum_{i}\kappa_{\lambda_{i}}}C_{\square\lambda_{2}\lambda^{T}_{3}}C_{\varnothing\lambda_{1}\lambda_{2}^{T}}C_{\varnothing\lambda_{3}\lambda_{1}^{T}}Tr_{\square}V,$
where the topological vertex is encoded by $C_{\lambda_{1},\lambda_{2},\lambda_{3}},$ a Laurent series in $q^{\frac{1}{2}}$ with integer coefficients indexed by three partitions. The closed topological string partition function $Z_{\mathbb{F}_{1}}$ of local $\mathbb{F}_{1}$ with exceptional curve $C$ is given by,
$Z_{\mathbb{F}_{1}}=\sum_{\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4}}(-1)^{\sum_{i}|\lambda_{i}|}Q^{(|\lambda_{4}|-|\lambda_{1}|-|\lambda_{3}|)C}Q^{(|\lambda_{1}|+|\lambda_{2}|+|\lambda_{3}|)H}q^{\sum_{i}\kappa_{\lambda_{i}}}C_{\varnothing\lambda_{1}\lambda_{4}^{T}}C_{\varnothing\lambda_{2}\lambda_{1}^{T}}C_{\varnothing\lambda_{3}\lambda_{2}^{T}}C_{\varnothing\lambda_{4}\lambda_{3}^{T}}$
[…]
Figure 5.2. The toric diagrams of local $\mathbb{C}\mathbb{P}^{2}$ (left) and local $\mathbb{F}_{1}$ (right). Internal edges have representations $R_{i}$ attached, where for the open calculation of local $\mathbb{C}\mathbb{P}^{2}$ with winding $1$, we have the fundamental representation $\square$ on the external edge.
We begin with degree one, $d=1$. Then all representation labels are either trivial or $\square$, simplifying the expressions for the vertex functions $C_{\lambda\mu\nu}$ (see pg.53, [AKMV] for explicit calculations). Extracting integer invariants, we get
(5.13) $n_{0}(K_{\mathbb{F}_{1}},\pi^{*}H-C)=-2,\quad n_{g}(K_{\mathbb{F}_{1}},\pi^{*}H-C)=0\text{ for }g>0$
and
(5.14) $n^{open}_{0}(K_{\mathbb{C}\mathbb{P}^{2}}/L,H+\beta_{0},(1))=2,\quad n^{open}_{g}(K_{\mathbb{C}\mathbb{P}^{2}}/L,H+\beta_{0},(1))=0\text{ for }g>0$
For degrees $2\leq d\leq 4$, we can use the calculations for the refined topological vertex in [IKV], Section 5.6.1, then un-refine them. These are listed in 1.
$d$ $\sum_{j_{L},j_{R}}N_{C}^{(j_{L},j_{R})}(j_{L},j_{R})$ $\widehat{f}_{\square\,\varnothing\,\varnothing}$
2 $(0,2)$ $5$
3 $(\frac{1}{2},4)\oplus(0,\frac{7}{2})\oplus(0,\frac{5}{2})$ $-(32+9z)$
4 $\left(\frac{3}{2},\frac{13}{2}\right)\oplus(1,6)\oplus(1,5)\oplus 2\left(\frac{1}{2},\frac{11}{2}\right)\oplus(0,6)\oplus 2\left(\frac{1}{2},\frac{9}{2}\right)$
$\oplus(0,5)\oplus\left(\frac{1}{2},\frac{7}{2}\right)\oplus 2(0,4)\oplus(0,3)\oplus(0,2)$ $286+288z+108z^{2}+14z^{3}$
Table 1. Open and closed BPS degeneracies.
The table lists the Gopakumar-Vafa invariants of local $\mathbb{F}_{1}$ in curve classes $(d-1)B+dF$, with the $SU(2)_{L}\times SU(2)_{R}$ spin content of BPS states supported on those curves. To un-refine, we want to ignore $j_{R},$ meaning each such representation contributes a factor of $(2j_{R}+1),$ so $(j_{L},j_{R})$ generates a term $(2j_{R}+1)(q^{-2j_{L}}+q^{-2j_{L}+2}+\ldots+q^{2j_{L}-2}+q^{2j_{L}})$ to the curve counting function (see also [HKR], Appendix B for some of these BPS invariants). For given $d$, we compute the total contribution from all of the $(j_{L},j_{R})$. After making a change of variables $z:=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2}$, the computed expression matches, up to a convention-dependent overall sign, the coefficient of $Q^{d}$ in $-\widehat{f}_{\square\,\varnothing\,\varnothing}$ (defined in Section 5.1) containing the open-BPS invariants in curve class $dH+\beta_{0}$. [^7] In Table 1, the coefficient of $(0,4)$ is listed in [IKV] as 1 instead of 2 — a typo, we believe. Indeed, the coefficient should be 2 in order to recover the genus-0 BPS invariant of local $\mathbb{F}_{1}$ in class $3B+4F$ in the limit $q\rightarrow 1$, which is 286. This invariant is computed in [LLW] [CKYZ] [AKMV]. This completes the verification of Corollary 5.4 for $\mathbb{C}\mathbb{P}^{2}$ for degrees $d\leq 4.$
Appendix A Combinatorial results
A.1. Laurent series inversion and substitution
Recall that we write $[x^{m}]f(x)$ to refer to the coefficient of $x^{m}$ in a series $f(x)$.
Let $f(x)$ be a formal Laurent series with complex coefficients of the form $f(x)=\sum_{k\geq-1}f_{k}x^{k}$ with $f_{-1}\neq 0$. A compositional inverse of $f$ is a series $g(z)$ that solves $f(g(z))=z$.
Lemma A.1 (Lagrange inversion for Laurent series).
As an element $g\in\mathbb{C}\llbracket z^{-1}\rrbracket$, there is a unique compositional inverse $g$ of $f$ given by the expression
$g(z)=\sum_{k>0}\frac{z^{-k}}{k}[x^{-1}]f^{k}.$
Moreover, $f(x)$ of the given form is uniquely determined by $g(z)$.
Proof.
The multiplicative inverse of $f$ can be written as
$\frac{1}{f(x)}=\frac{x}{h(x)},\quad h(x)=xf(x)=\sum_{k\geq-1}f_{k}x^{k+1}$
Note that $h(x)$ is a formal power series with $h(x)\neq 0$. The Lagrange inversion formula states that there exists a unique power series $\tilde{g}(z)$ which is the compositional inverse of $1/f(x)$. From this one can derive the Lagrange-Bürmann formula for formal power series, which gives an explicit expression for the coefficients of $\tilde{g}(z)$. This is [Sam], Theorem 7.5, which gives
$\tilde{g}(z)=\sum_{k>0}\frac{z^{k}}{k}[x^{k-1}]h^{k}=\sum_{k>0}\frac{z^{k}}{k}[x^{-1}]f^{k}.$
The equation $1/f(\tilde{g}(z))=z$ is equivalent to $f(\tilde{g}(z))=z^{-1}$ and to $f(\tilde{g}(z^{-1}))=z$. So the inverse of $f(x)$ is $g(z)=\tilde{g}(z^{-1})$ as claimed. Conversely, for any compositional inverse $g(z)$ of $f(x)$ we have that $\tilde{g}(z)=g(z^{-1})$ is a compositional inverse of $1/f(x)$. If $g(z)$ is a formal power series in $z^{-1}$, then $\tilde{g}(z)$ is a formal power series in $z$. This is unique by the Lagrange inversion formula for formal power series. Hence, $g(z)$ is unique as an element of $\mathbb{C}\llbracket z^{-1}\rrbracket$. The Lagrange inversion formula for formal power series also states that the power series $1/f(z)$ is uniquely determined by $\tilde{g}(z)$. Hence, $f(x)$ of the given form is uniquely determined by $g(z)$. ∎
Definition A.2.
The ring $\mathbb{C}\llbracket x^{-1}\rrbracket[x]$ is a field. An element $f\in\mathbb{C}\llbracket x^{-1}\rrbracket[x]$ takes the form
$f(x)=\sum_{k=-\infty}^{\operatorname{ord}(f)}f_{k}x^{k},\quad f_{\operatorname{ord}(f)}\neq 0,$
and we call $\operatorname{ord}(f)$ the order of $f$.
Lemma A.3.
Let
$f(x)=\sum_{k=-\infty}^{1}f_{k}x^{k}$
be a Laurent series with $\operatorname{ord}(f)=1$. There exists a unique and explicitly computable formal Laurent series $g(z)$ of order $\operatorname{ord}(g)=1$ that is the compositional inverse of $f$.
Proof.
By Lemma A.1 the series
$\tilde{f}(x):=f(x^{-1})=\sum_{k=-1}^{\infty}f_{-k}x^{k}$
has inverse
$\tilde{g}(z)=\sum_{k=1}^{\infty}\frac{z^{-k}}{k}[x^{-1}]\tilde{f}^{k}=\sum_{k=1}^{\infty}\frac{z^{-k}}{k}[x^{1}]f^{k}=f_{1}z^{-1}\left(1+\sum_{k=1}^{\infty}\frac{z^{-k}}{(k+1)f_{1}}[x^{1}]f^{k+1}\right).$
This means $\tilde{f}(\tilde{g}(z))=f(\tilde{g}(z)^{-1})=z$, hence $f(g(z))=z$ for
$g(z)=\tilde{g}(z)^{-1}=f_{1}^{-1}z\left(1+\sum_{k=1}^{\infty}\frac{z^{-k}}{(k+1)f_{1}}[x^{1}]f^{k+1}\right)^{-1},$
and this is a Laurent series with $\operatorname{ord}(g)=1$. ∎
Lemma A.4.
If $f\in\mathbb{C}\llbracket x^{-1}\rrbracket[x]$ and $f\neq 0$, then $f^{\prime}/f\in\mathbb{C}\llbracket x^{-1}\rrbracket\subset\mathbb{C}\llbracket x^{-1}\rrbracket[x]$ with $\operatorname{ord}(f^{\prime}/f)=-1$ and
$[x^{-1}](f^{\prime}/f)=\operatorname{ord}(f).$
Proof.
Write $n=\operatorname{ord}(f)$ and
$f(x)=\sum_{k=-\infty}^{n}f_{k}x^{k}.$
Then
$f^{\prime}(x)=\sum_{k=-\infty}^{n}kf_{k}x^{k-1}$
and
$\frac{f^{\prime}(x)}{f(x)}=\frac{nf_{n}x^{n-1}\left(1+\sum_{k=1}^{\infty}\frac{n-k}{n}\frac{f_{n-k}}{f_{n}}x^{-k}\right)}{f_{n}x^{n}\left(1+\sum_{k=1}^{\infty}\frac{f_{n-k}}{f_{n}}x^{-k}\right)}=nx^{-1}\left(1+\sum_{k=1}^{\infty}a_{k}x^{-1}\right)$
for some $a_{k}\in\mathbb{C}$. Hence, $\operatorname{ord}(f^{\prime}/f)=-1$ and $[x^{-1}](f^{\prime}/f)=n$ as claimed. ∎
Definition A.5.
For $f\in\mathbb{C}\llbracket x^{-1}\rrbracket[x]$, we write $\operatorname{Res}_{x}(f)=[x^{-1}]f$ and ${\sf Const}_{x}(f)=[x^{0}]f$. Moreover, for $f\in(\mathbb{C}\llbracket x^{-1}\rrbracket[x])\llbracket t\rrbracket$, we write $f=\sum_{k=0}^{\infty}f_{k}(x)t^{k}$ and define
$\operatorname{Res}_{x}(f)=\sum_{t=0}^{\infty}\operatorname{Res}_{x}(f_{k})t^{k},\qquad{\sf Const}_{x}(f)=\sum_{t=0}^{\infty}{\sf Const}_{x}(f_{k})t^{k}.$
There is a notion of formal derivation $f^{\prime}=\partial_{x}f$ for $f\in\mathbb{C}\llbracket x^{-1}\rrbracket[x]$ and then we define $\partial_{x}$ by $\mathbb{C}$-linear extension also for $f\in(\mathbb{C}\llbracket x^{-1}\rrbracket[x])\llbracket t\rrbracket$.
Lemma A.6.
Let $f\in(\mathbb{C}\llbracket x^{-1}\rrbracket[x])\llbracket t\rrbracket$ and $x\in\mathbb{C}\llbracket u^{-1}\rrbracket[u]\setminus\{0\}$. Then
$\operatorname{Res}_{u}\Big{(}f(x(u))x^{\prime}(u)\Big{)}=\operatorname{ord}(x)\operatorname{Res}_{x}\left(f\right),$
and
${\sf Const}_{u}\left(f(x(u))\frac{u\partial_{u}x(u)}{x(u)}\right)=\operatorname{ord}(x){\sf Const}_{x}(f).$
Proof.
Note that $\operatorname{im}(\partial_{x})=\operatorname{ker}(\operatorname{Res}_{x})$, so we can write $f(x)=\operatorname{Res}_{x}(f)x^{-1}+\partial_{x}\widetilde{f}(x)$. Insertion yields
$f(x(u))x^{\prime}(u)=\left(\operatorname{Res}_{x}(f)x(u)^{-1}+\partial_{x}\widetilde{f}(x(u))\right)x^{\prime}(u)=\operatorname{Res}_{x}(f)x(u)^{-1}x^{\prime}(u)+\partial_{u}\widetilde{f}(x(u)).$
Lemma A.4 gives $\operatorname{Res}_{u}(x^{\prime}(u)/x(u))=\operatorname{ord}(x)$ and $\operatorname{Res}_{u}(\partial_{u}\widetilde{f}(x(u))=0$, since $\operatorname{im}(\partial_{u})=\operatorname{ker}(\operatorname{Res}_{u})$. Hence, $\operatorname{Res}_{u}\left(f(x(u))x^{\prime}(u)\right)=\operatorname{ord}(x)\operatorname{Res}_{x}\left(f(x)\right)$ as claimed. The second statement follows from the first statement applied to $f(x)/x$. ∎
A.2. Bell polynomial identities
In the following we consider polynomials and power series over a $\mathbb{C}$-algebra $A$ that is an integral domain.
Definition A.7.
For a sequence of complex numbers $a=(a_{1},a_{2},\ldots)$ and $k\geq 0$, let $\widehat{B}_{n,k}(a)$ be the coefficient of $x^{n}$ in $(a_{1}x+a_{2}x^{2}+\ldots)^{k}$,
$\left(\sum_{i=1}^{\infty}a_{i}x^{i}\right)^{k}=\sum_{n=1}^{\infty}\widehat{B}_{n,k}(a)x^{n}.$
For $n<k$ we have $\widehat{B}_{n,k}(a)=0$ and we define $\widehat{B}_{n,k}(a):=0$ whenever $k<0$.
Proposition A.8.
$\widehat{B}_{n,k}(a)$ is equal to the partial ordinary Bell polynomial
$\widehat{B}_{n,k}(a)=\widehat{B}_{n,k}(a_{1},\ldots,a_{n-k+1})=\sum_{\begin{subarray}{c}(j_{1},\ldots,j_{n-k+1})\in\mathbb{Z}_{\geq 0}^{n-k+1}\\ \sum j_{i}=k\\ \sum ij_{i}=n\end{subarray}}\frac{k!}{j_{1}!\cdots j_{n-k+1}!}a_{1}^{j_{1}}\cdots a_{n-k+1}^{j_{n-k+1}}.$
Proof.
E.T. Bell introduced the Bell polynomials for this property, see [Bel]. ∎
Recall that we write $[x^{m}]g(x)$ to refer to the coefficient of $x^{m}$ in $g(x)$.
Lemma A.9.
For a power series $f=\sum_{k\geq 0}a_{k}x^{k}$ and $n,m\geq 0$, we have
$[x^{m}]f^{n}=\sum_{k=0}^{m}\binom{n}{k}a_{0}^{n-k}\widehat{B}_{m,k}(a).$
Proof.
For $a_{0}=0$ the claimed equation reduces to $[x^{m}]f^{n}=\widehat{B}_{m,n}(a)$, which is Proposition A.8. In general, we can write $f(x)=a_{0}+g(x)$, where $g(x)=\sum_{k>0}a_{k}x^{k}$ has no constant term. Then
$[x^{m}]f^{n}=\sum_{k=0}^{m}\binom{n}{k}a_{0}^{n-k}[x^{m}]g^{k}.$
It remains to observe that $[x^{m}]g^{k}=\widehat{B}_{m,k}(a)$ since $g$ has no constant term and otherwise the same coefficients as $f$. This completes the proof. ∎
Lemma A.10.
Let $n,k\in\mathbb{Z}_{\geq 0}$, $\tau\in\mathbb{C}$, and $a=(a_{1},a_{2},\ldots)$. Let $\alpha(l,m)$ be a polynomial in $l$ and $m$ of degree at most one taking non-zero values for $l,m\in\{0,...,n\}$. We have
$\binom{\tau}{k}\widehat{B}_{n,k}(a)=\sum_{m=0}^{n}\sum_{l=0}^{m}\frac{\alpha(0,0)}{\alpha(l,m)}\binom{\tau-\alpha(l,m)}{k-l}\binom{\alpha(l,m)}{l}\widehat{B}_{m,l}(a)\widehat{B}_{n-m,k-l}(a).$
Proof.
The exponential Bell polynomials $B_{n,k}(a)$ are by definition related to $\widehat{B}_{n,k}(a)$ by
(A.1) $\widehat{B}_{n,k}(a_{1},\ldots,a_{n-k+1})=\frac{k!}{n!}B_{n,k}(1!a_{1},\ldots,(n-k+1)!a_{n-k+1}).$
The second equation of [BGW], Corollary 10, is
$\binom{\tau}{k}B_{n,k}(x)=\sum_{l=0}^{k}\sum_{m=l}^{n}\frac{\alpha(0,0)}{\alpha(l,m)}\binom{\tau-\alpha(l,m)}{k-l}\binom{\alpha(l,m)}{l}\frac{\binom{n}{m}}{\binom{k}{l}}B_{m,l}(x)B_{n-m,k-l}(x).$
We substitute $x_{i}=i!a_{i}$ and use (A.1) to find that $\binom{\tau}{k}B_{n,k}(a)$ equals
$\sum_{l=0}^{k}\sum_{m=l}^{n}\frac{\alpha(0,0)}{\alpha(l,m)}\binom{\tau-\alpha(l,m)}{k-l}\binom{\alpha(l,m)}{l}\underbrace{\frac{\binom{n}{m}}{\binom{k}{l}}\frac{k!}{n!}\frac{m!}{l!}\frac{(n-m)!}{(k-l)!}}_{=1}B_{m,l}(a)B_{n-m,k-l}(a).$
The binomials and factorials in the last sum that do not show in the asserted identity combine to the factor $1$. It remains to observe that the last summation only differs from the one in the assertion by removing terms where $l>k$ and these terms contain a zero factor $\widehat{B}_{n-m,k-l}(a)$, so we have verified the assertion. ∎
Lemma A.11.
For a power series $f(x)=1+\sum_{k\geq 1}f_{k}x^{k}$ one has, for all $n\geq 0$,
$[x^{n-1}]f^{n}=\sum_{\begin{subarray}{c}a>0,b\geq 0\\ a+b=n\end{subarray}}\frac{1}{a}[x^{a-1}]f^{a}[x^{b}]f^{b}.$
Proof.
By Lemma A.9, the right hand side of the claim can be written as
$\sum_{\begin{subarray}{c}a>0,b\geq 0\\ a+b=n\end{subarray}}\frac{1}{a}\left(\sum_{k_{1}=0}^{a-1}\binom{a}{k_{1}}\widehat{B}_{a-1,k_{1}}(f)\right)\left(\sum_{k_{2}=0}^{b}\binom{b}{k_{2}}\widehat{B}_{b,k_{2}}(f)\right)$
$=$ $\sum_{\begin{subarray}{c}a>0,b\geq 0\\ a+b=n\end{subarray}}\sum_{k_{1}=0}^{a-1}\sum_{k_{2}=0}^{b}\frac{1}{a}\binom{a}{k_{1}}\binom{b}{k_{2}}\widehat{B}_{a-1,k_{1}}(f)\widehat{B}_{b,k_{2}}(f)$
=$\stackrel{{\scriptstyle\begin{subarray}{c}k:=k_{1}+k_{2}\\ b=n-a\end{subarray}}}{{=}}$ $\sum_{k=0}^{n-1}\sum_{a=1}^{n}\sum_{k_{1}=0}^{a-1}\frac{1}{a}\binom{a}{k_{1}}\binom{n-a}{k-k_{1}}\widehat{B}_{a-1,k_{1}}(f)\widehat{B}_{n-a,k-k_{1}}(f)$
=$\stackrel{{\scriptstyle\begin{subarray}{c}m:=a-1\\ l:=k_{1}\end{subarray}}}{{=}}$ $\sum_{k=0}^{n-1}\sum_{m=0}^{n-1}\sum_{l=0}^{m}\frac{1}{m+1}\binom{m+1}{l}\binom{n-(m+1)}{k-l}\widehat{B}_{m,l}(f)\widehat{B}_{n-m-1,k-l}(f)$
[…]
In the last equality we used Lemma A.9 and in the one before we used Lemma A.10 for the parameters $n-1,k,\tau=n$ and $\alpha(l,m)=m+1$. ∎
Proposition A.12.
For a power series $f(x)=1+\sum_{k\geq 1}f_{k}x^{k}$ one has
$\textup{exp}\left(\sum_{k>0}\frac{1}{k}[x^{k}]f^{k}z^{k}\right)=\sum_{k>0}\frac{1}{k}[x^{k-1}]f^{k}z^{k-1}.$
Proof.
These functions have the same value $1$ at $z=0$, so it is enough to show that their derivatives agree [^8] We thank Ezra Getzler for suggesting this approach.. Deriving the equation $\textup{exp}(A)=B$ gives $BA^{\prime}=B^{\prime}$, thus it suffices to show
$\left(\sum_{k>0}\frac{1}{k}[x^{k-1}]f^{k}z^{k-1}\right)\left(\sum_{k>0}[x^{k}]f^{k}z^{k-1}\right)=\sum_{k>0}\frac{k-1}{k}[x^{k-1}]f^{k}z^{k-2}.$
By comparing the $z^{n-2}$-coefficients, this translates, for every $n$, into the identity
$\sum_{\begin{subarray}{c}a,b>0\\ a+b=n\end{subarray}}\frac{1}{a}[x^{a-1}]f^{a}[x^{b}]f^{b}=\frac{n-1}{n}[x^{n-1}]f^{n}.$
We can add $\frac{1}{n}[x^{n-1}]f^{n}$ to both sides and incorporate it into the sum as $b=0$ to get the equivalent identity
$\sum_{\begin{subarray}{c}a>0,b\geq 0\\ a+b=n\end{subarray}}\frac{1}{a}[x^{a-1}]f^{a}[x^{b}]f^{b}=[x^{n-1}]f^{n}.$
that we recognize as the statement of Lemma A.11 which concludes the proof. ∎